### v2

# Unique Ergodicity of Stochastic Theta Method for Monotone SDEs driven by Nondegenerate Multiplicative Noise

Abstract.
We establish the unique ergodicity of the Markov chain generated by the stochastic theta method (STM) with $\theta\in[1/2,1]$ for monotone SODEs, without growth restriction on the coefficients, driven by nondegenerate multiplicative noise. The main ingredient of the arguments lies in constructing new Lyapunov functions involving the coefficients, the stepsize, and $\theta$, and the irreducibility and the strong Feller property for the STM. We also generalize the arguments to the temporal drift-implicit Euler (DIE) method and its Galerkin-based full discretizations for a class of monotone SPDEs driven by infinite-dimensional nondegenerate multiplicative trace-class noise. Applying these results to the stochastic Allen–Cahn equation indicates that its DIE scheme is uniquely ergodic for any interface thickness, which gives an affirmative answer to a question proposed in (J. Cui, J. Hong, and L. Sun, Stochastic Process. Appl. (2021): 55–93). Numerical experiments verify our theoretical results.
Key words and phrases:
monotone stochastic differential equation, numerical invariant measure, numerical ergodicity, stochastic Allen–Cahn equation, Lyapunov structure
2010 Mathematics Subject Classification:
Primary 60H35; 60H15, 65M60
The first author is supported by the National Natural Science Foundation of China, No. 12101296, Shenzhen Basic Research Special Project (Natural Science Foundation) Basic Research (General Project), No. JCYJ20220530112814033, and Basic and Applied Basic Research Foundation of Guangdong Province, No. 2024A1515012348.
1. Introduction
The long-time behavior of Markov processes generated by stochastic differential equations (SDEs) is a natural and intriguing question and has been investigated in recent decades. As a significant long-time behavior, the ergodicity characterizes the case of temporal average coinciding with spatial average, which has a lot of applications in quantum mechanics, fluid dynamics, financial mathematics, and many other fields [12, 17]. The spatial average, i.e., the mean of a given test function for the invariant measure of the considered Markov process, also known as the ergodic limit, is desirable to compute in practical applications. One has to investigate a stochastic system over long time intervals, which is one of the main difficulties from the computational perspective. It is well-known that the explicit expression of the invariant measure for a stochastic nonlinear system is rarely available; exceptional examples are gradient Langevin systems driven by additive noise, e.g., [3, 20]. For this reason, it motivated and fascinated a lot of investigations in the recent decade for constructing numerical algorithms that can inherit the ergodicity of the original system and approximate the ergodic limit efficiently.
Much progress has been made in the design and analysis of numerical approximations of the desired ergodic limits for SDEs under a strong dissipative condition so that the Markov chains are contractive, see, e.g., [22, 34] and references therein for numerical ergodicity of backward Euler or Milstein schemes for dissipative SODEs, [21] for dissipative SODEs with Markovian switching, and [4, 5, 7, 8, 9, 11, 24] for approximating the invariant measures via temporal tamed, Galerkin-based linearity-implicit Euler or exponential Euler schemes, and high order integrator for parabolic SPDEs driven by additive noise. See also [28] and [15] for numerical ergodicity of backward Euler scheme and its versions for monotone SODEs and spectral Galerkin approximation for 2-D stochastic Navier–Stokes equations, respectively, both driven by additive degenerate noise.
We note that most of the above works of literature focus on the numerical ergodicity of strong dissipative SDEs driven by additive noise; the numerical ergodicity for weak dissipative SDEs in the multiplicative noise case is more subtle. On the other hand, the authors in [11] investigated the temporal DIE and the spectral Galerkin scheme to approximate the invariant measure of the stochastic Allen–Cahn equation (SACE) and proposed a question of whether the invariant measure of the DIE scheme is unique. It should be pointed out that, starting from a fixed state space, the corresponding Markov chains generated by numerical schemes of a class of SPDEs, including the SACE, always lie in its proper subspace. This difference from the finite-dimensional case makes the numerical ergodicity of SPDEs more challenging.
The above two questions on the unique ergodicity of numerical approximations for monotone SDEs in both finite-dimensional and infinite-dimensional settings motivate the present study. Our main aim is to establish the unique ergodicity of the STM (see (STM) with $\theta\in[1/2,1]$), including the numerical schemes studied in [16, 28, 31] for monotone SODEs, without growth restriction on the coefficients, driven by nondegenerate multiplicative noise. The main ingredient of our arguments lies in constructing a new Lyapunov function, combined with the irreducibility and strong Feller property for the STM.
It is not difficult to show that $V(\cdot)=\left\lvert\cdot\right\rvert^{2}$ is a natural Lyapunov function of the considered monotone (SDE). However, it was shown in [18] that the Euler–Maruyama scheme (i.e., (STM) with $\theta=0$) applied to (SDE) with superlinear growth coefficients would blow up in $p$-th moment for all $p\geq 2$. Therefore, $\left\lvert\cdot\right\rvert^{2}$ is not an appropriate Lyapunov function of this scheme in the setting of the present study, and we mainly focus on the case $\theta\in[1/2,1]$. By exploring the monotone structure of the coefficients and martingale property of the driven Wiener process, we construct several Lyapunov functions (see, e.g., (3.2) and (3.6)) for the STM, which involve both the coefficients, the stepsize, and $\theta$.
Then we generalize our methodology to the temporal DIE scheme (see (DIE)) and its Galerkin-based fully discrete schemes (see (DIEG)), which have been studied in [11, 14, 19, 25, 27], for monotone SPDEs with polynomial growth coefficients driven by infinite-dimensional nondegenerate multiplicative trace-class noise. Apart from the Lyapunov structure, the main ingredient in this part is the rigorous proof of the irreducibility and strong Feller property in the underlying Hilbert space (say $H$), even though the corresponding Markov chain always lies in a smaller Hilbert space (say $\dot{H}^{1}$); see Proposition 4.1. Applying these results to SACE driven by nondegenerate noise indicates that the DIE scheme and its Galerkin-based full discretizations are uniquely ergodic, respectively, for any interface thickness, which gives an affirmative answer to the question proposed in [11].
The paper is organized as follows. In Section 2, we give the main assumptions on monotone SODEs and recall the ergodic theory of Markov chains that will be used throughout. The Lyapunov structure of the STM, with irreducibility and strong Feller property, is explored in Section 3. In Section 4, we generalize the arguments in Section 3 to monotone SPDEs, including the SACE. The theoretical results are validated by numerical experiments in Section 5. We include some discussions in the last section.
2. Preliminaries
In this section, we present the required main assumptions and recall the general ergodic theory of Markov chains that will be used throughout the paper.
2.1. Main Assumptions
Let us first consider the $d$-D SODE
(SDE) $\mathrm{d}X(t)=b(X(t))\mathrm{d}t+\sigma(X(t))\mathrm{d}W(t),~{}t\geq 0,$
driven by an $\mathbb{R}^{m}$-valued Wiener process $W$ on a complete filtered probability space $(\Omega,\mathscr{F},\mathbb{F}:=(\mathscr{F}(t))_{t\geq 0},\mathbb{P})$, where $b:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $\sigma:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times m}$ are measurable functions.
Our main focus is on the invariant measure of the Markov chain generated by the following stochastic theta method (STM) with $\theta\in[0,1]$:
(STM) $X_{j+1}=X_{j}+(1-\theta)b(X_{j})\tau+\theta b(X_{j+1})\tau+\sigma(X_{j})\delta_{j}W,$
where $\tau\in(0,1)$ is a fixed step-size, $\delta_{j}W:=W(t_{j+1})-W(t_{j})$ with $t_{j}=j\tau$, $j\in\mathbb{N}$. When $\theta=0,1/2,1$, it is called the Euler–Maruyama scheme, the mid-point scheme, and the backward Euler method, respectively.
[…]
Assumption 2.1.
There exists a constant $L_{1}\in\mathbb{R}$ such that
(2.1) $2\left\langle b(x)-b(y),x-y\right\rangle+\left\lVert\sigma(x)-\sigma(y)\right\rVert^{2}\leq L_{1}\left\lvert x-y\right\rvert^{2},\quad\forall~{}x,y\in\mathbb{R}^{d}.$
Under the monotone condition (2.1), one can show the existence and uniqueness of the $(\mathscr{F}_{t})_{t\geq 0}$-adapted solution to (SDE), see, e.g., [23, Theorem 3.1.1]. The following result gives the unique solvability of (STM).
Lemma 2.1.
Let Assumption 2.1 hold. Then (STM) can be uniquely solved when $L_{1}\theta\tau<2$.
Proof.
Define
(2.2) $\hat{b}(x)=x-\theta b(x)\tau,\quad x\in\mathbb{R}^{d}.$
Then (STM) becomes
$\hat{b}(X_{j+1})=X_{j}+(1-\theta)b(X_{j}^{h})\tau+\sigma(X_{j})\delta_{j}W,\quad j\in\mathbb{N}.$
It follows from (2.1) that
$\langle x-y,\hat{b}(x)-\hat{b}(y)\rangle$ $=\left\langle x-y,x-y-(b(x)-b(y))\right\rangle\theta\tau$
$=\left\lVert x-y\right\rVert^{2}-\left\langle x-y,b(x)-b(y)\right\rangle{\color[rgb]{1,.5,0}\theta}\tau$
$\geq(1-L_{1}\theta\tau/2)\left\lVert x-y\right\rVert^{2}.$
As $L_{1}\theta\tau<2$, $\hat{b}$ defined in (2.2) is uniformly monotone so that it is invertible (see, e.g., [33, Theorem C.2]), from which we can conclude the unique solvability of (STM). ∎
Assumption 2.2.
There exist two positive constants $L_{2}$ and $L_{3}$ such that
(2.3) $2\left\langle b(x),x\right\rangle+\left\lVert\sigma(x)\right\rVert^{2}\leq L_{2}-L_{3}\left\lvert x\right\rvert^{2},\quad\forall~{}x\in\mathbb{R}^{d}.$
We also need the following continuity of $b$ and non-degeneracy of $\sigma$.
Assumption 2.3.
$b$ is continuously differentiable and for any $x\in\mathbb{R}^{d}$, $\sigma(x)\sigma(x)^{\top}$ is positive definite on $\mathbb{R}^{m}$ with inverse $[\sigma\sigma^{\top}]^{-1}$.
It is clear from Assumption 2.3 that the determinant $\det J_{b}(y)$ of the Jacobi matrix $J_{b}$ of $b$ is continuous, which will be needed in the proof of Corollary 3.2.
Example 2.1.
A motivating example for $b$ and $\sigma$ satisfying Assumptions 2.1, 2.2, and 2.3 is
$b(x)=x-x^{3},\quad\sigma(x)=\sqrt{x^{2}+1},\quad x\in\mathbb{R}.$
Clearly, Assumption 2.3 holds. As $\sigma$ is Lipschitz continuous with Lipschitz constant 1, we have
$\left\langle b(x)-b(y),x-y\right\rangle+\|\sigma(x)-\sigma(y)\|^{2}$
$=(x-y)^{2}(1-x^{2}-xy-y^{2})+|\sqrt{x^{2}+1}-\sqrt{y^{2}+1}|$
$\leq\left\lvert x-y\right\rvert^{2}(2-x^{2}-xy-y^{2})\leq 2\left\lvert x-y\right\rvert^{2},$
$2\left\langle b(x),x\right\rangle+\|\sigma(x)\|^{2}=2x^{2}-2x^{4}+x^{2}+1\leq 3-x^{2}.$
as $x^{2}+xy+y^{2}\geq 0$, for all $x,y\in\mathbb{R}$. This shows Assumptions 2.1 and 2.2 with $L_{1}=2$ and $L_{2}=3$, $L_{3}=1$, respectively.
Under Assumptions 2.1, 2.2, and 2.3, (SDE) is ergodic, see, e.g., [35] for the use of coupling methods based on a changing measure technique in combination of Girsanov theorem. In the discrete-time model, such a Girsanov theorem does not exist. Instead, we will utilize the general ergodic theory of Markov chains.
2.2. Preliminaries on Ergodicity of Markov Chains
Let $(H,\|\cdot\|)$ (either $\mathbb{R}^{d}$ in Section 3 and $L^{2}$-space in Section 4) be a separable Hilbert space equipped with Borel $\sigma$-algebra $\mathcal{B}(H)$. Let $(X_{n})_{n\in\mathbb{N}}$ be an $H$-valued Markov chain with transition kernel $P(x,B)$, $x\in H$, $B\in\mathcal{B}(H)$.
Recall that a set $C\in\mathcal{B}(H)$ is called an ($(m,\nu)$-)small set if there exists an integer $m>0$ and a non-trivial measure $\nu$ on $\mathcal{B}(H)$ such that
$P^{m}(x,B)\geq\nu(B),\quad\forall~{}x\in C,~{}B\in\mathcal{B}(H).$
$(X_{n})_{n\in\mathbb{N}}$ is said to satisfy the minorization condition (with $C$) if there exists some $\delta>0$, $C\in\mathcal{B}(H)$, and probability measure $\nu$ with $\nu(C)=1$ such that
(2.4) $P(x,B)\geq\delta\chi_{C}(x)\nu(B),\quad\forall~{}x\in H,~{}B\in\mathcal{B}(H).$
It is clear that if $(X_{n})_{n\in\mathbb{N}}$ satisfies the minorization condition (2.4) with $C$, then $C$ is a $(1,\delta\nu)$-small set. $(X_{n})_{n\in\mathbb{N}}$ is called (open set) irreducible if $P^{n}(x,A)>0$ for any $x\in H$, any non-empty open set $A$ in $H$ and some $n\in\mathbb{N}_{+}$; it is called strong Feller if $P(\cdot,A)$ is lower continuous for any Borel set $A\in\mathcal{B}(H)$. We refer to [29] for more details on Markov chains.
A probability measure $\mu$ in $H$ is called invariant for the Markov chain $(X_{n})_{n\in\mathbb{N}}$ or its transition kernel $P$, if
$\int_{H}P\phi(x)\mu({\rm d}x)=\mu(\phi):=\int_{H}\phi(x)\mu({\rm d}x),\quad\forall~{}\phi\in\mathcal{C}_{b}(H).$
This is equivalent to $\int_{H}P(x,A)\mu({\rm d}x)=\mu(A)$ for all $A\in\mathcal{B}(H).$ An invariant (probability) measure $\mu$ is called ergodic for $(X_{n})_{n\in\mathbb{N}}$ or $P$, if
(2.5) $\lim_{m\to\infty}\frac{1}{m}\sum_{k=0}^{m}P_{k}\phi(x)=\mu(\phi)\quad\text{in}~{}L^{2}(H;\mu),\quad\forall~{}\phi\in L^{2}(H;\mu).$
It is well-known that if $(X_{n})_{n\in\mathbb{N}}$ admits a unique invariant measure, then it is ergodic; in this case, we call it uniquely ergodic.
The unique ergodicity for a Markov chain is equivalent to the one-step Lyapunov condition
(2.6) $PV(x)-V(x)\leq-1+c\chi_{C}(x),\quad\forall~{}x\in H,$
for some $(m,\nu)$-small set $C$, constant $c$, and function $V:H\to[0,\infty)$ (called a Lyapunov function) which is finite $\nu$-almost everywhere (see, e.g., [29, Theorem 1.3.1]). A stronger but more adapted Lyapunov condition is that there exists $\rho\in(0,1)$ and $\kappa\geq 0$ such that
(2.7) $PV\leq\rho V+\kappa.$
Indeed, if (2.7) holds, then (2.6) holds with $C=\{V\leq(\kappa+1)/(1-\rho)\}$ and a large $c$. Under the stronger Lyapunov condition (2.7), one could obtain the geometric ergodicity if the minorization condition (2.4) holds for $C=\{x:V(x)\leq 2\kappa/(\gamma-\rho)\}$ with some $\gamma\in(\rho^{1/2},1)$ (see [28, Theorem 2.5]).
3. Ergodicity of STM for Monotone SODEs
Denote by $\mathcal{A}$ the infinitesimal generator of (SDE), .e.,
$\mathcal{A}:=\sum_{i=1}^{d}b^{i}\frac{\partial}{\partial x_{i}}+\frac{1}{2}\sum_{i,j=1}^{d}(\sigma\sigma^{\top})_{ij}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}.$
Under Assumption 2.2, it is not difficult to show that
$\mathcal{A}\left\lvert X(t)\right\rvert^{2}=2\left\langle b(X(t)),X(t)\right\rangle+\left\lVert\sigma(X(t))\right\rVert^{2}\leq L_{2}-L_{3}\left\lvert X(t)\right\rvert^{2},\quad t\geq 0.$
Hence, $V(\cdot)=\left\lvert\cdot\right\rvert^{2}$ is a natural Lyapunov function of (SDE). However, it was shown in [18] that for (STM) with $\theta=0$ applied to (SDE) with superlinear growth coefficients, $\lim_{n\to\infty}\mathbb{E}|X_{n}|^{p}=\infty$ for all $p\geq 2$ so that $V(x)=\left\lvert x\right\rvert^{2}$ is not an appropriate Lyapunov function of this scheme.
In this section, we first construct a Lyapunov function and then derive the irreducibility and existence of continuous density for (STM) with $\theta\in[1/2,1]$.
3.1. Lyapunov structure
We begin with the following elementary inequality, which will be helpful to explore the Lyapunov structure of (STM). It is a straightforward generalization of [31, Lemma 3.3].
Lemma 3.1.
Let Assumption 2.2 hold. Then for any $\rho,\beta\in\mathbb{R}$ with $\rho\geq\beta\geq 0$, there exists $C=\frac{1+\beta L_{3}}{1+\rho L_{3}}$ such that
(3.1) $\left\lvert x-\beta b(x)\right\rvert^{2}+\beta L_{2}+(C\rho-\beta)\|\sigma(x)\|^{2}\leq C(\left\lvert x-\rho b(x)\right\rvert^{2}+\rho L_{2}),\quad x\in\mathbb{R}^{d}.$
Proof.
Let $x\in\mathbb{R}^{d}$. Direct estimation yields that
$\left\lvert x-\beta b(x)\right\rvert^{2}+\beta L_{2}+(C\rho-\beta)\|\sigma(x)\|^{2}-C(\left\lvert x-\rho b(x)\right\rvert^{2}+\rho L_{2})$
$=$ $(1-C)\left\lvert x\right\rvert^{2}+(C\rho-\beta)[2\left\langle x,b(x)\right\rangle-L_{2}+\|\sigma(x)\|^{2}]+(\beta^{2}-C\rho^{2})\left\lvert b(x)\right\rvert^{2}.$
It follows from (2.3) that the above equality can be bounded by
$[(1-C)-(C\rho-\beta)L_{3}]\left\lvert x\right\rvert^{2}+(\beta^{2}-C\rho^{2})\left\lvert b(x)\right\rvert^{2},$
from which we conclude (3.1) by noting that $1-C=(C\rho-\beta)L_{3}$ and $\beta^{2}-C\rho^{2}\leq 0$ with $C=\frac{1+\beta L_{3}}{1+\rho L_{3}}$. ∎
We have the following strong Lyapunov structure for (STM) with $\theta\in(1/2,1]$ under the coupled monotone and coercive conditions (2.1) and (2.3), respectively, which is crucial to derive the geometric ergodicity of (STM) with $\theta\in(1/2,1]$ in the next part.
Theorem 3.1.
Let Assumption 2.1 and 2.2 hold. Then for any $\theta\in(1/2,1]$, $\tau\in(0,1)$ with $L_{1}\theta\tau<2$, and $\lambda\in(0,2\theta-1]$, there exists a Lyapunov function $V_{\theta}:\mathbb{R}^{d}\to[0,\infty)$ defined by
(3.2) $V_{\theta}(x)=\left\lvert x-(1-\theta+\lambda)b(x)\tau\right\rvert^{2}+(2\theta-1-\lambda)\left\lVert\sigma(x)\right\rVert^{2}\tau,\quad x\in\mathbb{R}^{d},$
such that (2.7) holds for (STM) with $\rho=\frac{1+(1-\theta)L_{3}\tau}{1+(1-\theta+\lambda)L_{3}\tau}\in(0,1)$ and $\kappa=L_{2}\tau>0$, i.e.,
(3.3) $\mathbb{E}[V_{\theta}(X_{n+1})\mid\mathscr{F}_{n}]\leq\frac{1+(1-\theta)L_{3}\tau}{1+(1-\theta+\lambda)L_{3}\tau}V(X_{n})+L_{2}\tau,\quad n\in\mathbb{N}.$
Proof.
For $n\in\mathbb{N}$, to simplify the notation, denote by $b_{n}:=b(X_{n})$ and $\sigma_{n}:=\sigma(X_{n})$, and define
$F_{n}:=\left\lvert X_{n}-(1-\theta+\lambda)b_{n}\tau\right\rvert^{2}+L_{2}(1-\theta+\lambda)\tau.$
It is not difficult to show that
$F_{n+1}=$ $\left\lvert X_{n+1}-\theta b_{n+1}\tau+(2\theta-1-\lambda)b_{n+1}\tau\right\rvert^{2}+L_{2}(1-\theta+\lambda)\tau$
$=$ $\left\lvert X_{n+1}-\theta b_{n+1}\tau\right\rvert^{2}+2(2\theta-1-\lambda)[\left\langle X_{n+1},b_{n+1}\right\rangle\tau-\theta\left\lvert b_{n+1}\right\rvert^{2}\tau^{2}]$
$+(2\theta-1-\lambda)^{2}\left\lvert b_{n+1}\right\rvert^{2}\tau^{2}+L_{2}(1-\theta+\lambda)\tau$
$=$ $\left\lvert X_{n+1}-\theta b_{n+1}\tau\right\rvert^{2}+2(2\theta-1-\lambda)\left\langle X_{n+1},b_{n+1}\right\rangle\tau$
$+(\lambda+1-2\theta)(\lambda+1)\left\lvert b_{n+1}\right\rvert^{2}\tau^{2}+L_{2}(1-\theta+\lambda)\tau.$
Substituting (STM) in the first term on the right-hand side of the above equality, we obtain
$\left\lvert X_{n+1}-\theta b_{n+1}\tau\right\rvert^{2}=$ $\left\lvert X_{n}+(1-\theta)b_{n}\tau+\sigma_{n}\delta_{n}W\right\rvert^{2}$
$=$ $\left\lvert X_{n}+(1-\theta)b_{n}\tau\right\rvert^{2}+2\left\langle X_{n}+(1-\theta)b_{n}\tau,\sigma_{n}\delta_{n}W\right\rangle+\left\lvert\sigma_{n}\delta_{n}W\right\rvert^{2}$
$=$ $\left\lvert X_{n}-(1-\theta)b_{n}\tau\right\rvert^{2}+4(1-\theta)\left\langle X_{n},b_{n}\right\rangle\tau$
$+2\left\langle X_{n}+(1-\theta)b_{n}\tau,\sigma_{n}\delta_{n}W\right\rangle+\left\lvert\sigma_{n}\delta_{n}W\right\rvert^{2}.$
Then we have
(3.4) $\begin{split}F_{n+1}=&\left\lvert X_{n}-(1-\theta)b_{n}\tau\right\rvert^{2}+4(1-\theta)\left\langle X_{n},b_{n}\right\rangle\tau+2(2\theta-1-\lambda)\left\langle X_{n+1},b_{n+1}\right\rangle\tau\\ &+(\lambda+1-2\theta)(\lambda+1)\left\lvert b_{n+1}\right\rvert^{2}\tau^{2}+\left\lVert\sigma_{n}\right\rVert^{2}\tau+L_{2}(1-\theta+\lambda)h+O_{n},\end{split}$
where
$O_{n}=2\left\langle X_{n}+(1-\theta)b_{n}\tau,\sigma_{n}\delta_{n}W\right\rangle+\left\lvert\sigma_{n}\delta_{n}W\right\rvert^{2}-\left\lVert\sigma_{n}\right\rVert^{2}\tau.$
It is not difficult to show that $\mathbb{E}[O_{n}\mid\mathscr{F}_{n}]=0$ by using properties of conditional expectation and the Wiener process.
Following (3.4) and Lemma 3.1 with $\rho=(1-\theta+\lambda)\tau$, $\beta=(1-\theta)\tau$, $C=N_{\tau}:=\frac{1+(1-\theta)L_{3}\tau}{1+(1-\theta+\lambda)L_{3}\tau}\in(0,1)$, and $C\rho-\beta=B_{\tau}:=\frac{\lambda\tau}{1+(1-\theta+\lambda)L_{3}\tau}$, and discarding the term $(\lambda+1-2\theta)(\lambda+1)\left\lvert b_{n+1}\right\rvert^{2}\tau^{2}$ under the assumption $10<\lambda\leq 2\theta-$ 1, we have
$F_{n+1}\leq$ $\left\lvert X_{n}-(1-\theta)b_{n}\tau\right\rvert^{2}+(1-\theta)L_{2}\tau+4(1-\theta)\left\langle X_{n},b_{n}\right\rangle\tau+\left\lVert\sigma_{n}\right\rVert^{2}\tau$
$+2(2\theta-1-\lambda)\left\langle X_{n+1},b_{n+1}\right\rangle\tau+L_{2}\lambda\tau+O_{n}$
$\leq$ $N_{\tau}F_{n}-B_{\tau}\|\sigma_{n}\|^{2}\tau+4(1-\theta)\left\langle X_{n},b_{n}\right\rangle\tau+\left\lVert\sigma_{n}\right\rVert^{2}\tau$
$+2(2\theta-1-\lambda)\left\langle X_{n+1},b_{n+1}\right\rangle\tau+L_{2}\lambda\tau+O_{n}$
$\leq$ $N_{\tau}F_{n}+2(1-\theta)(L_{2}-L_{3}|X_{n}|^{2}-\left\lVert\sigma_{n}\right\rVert^{2})\tau+(1-B_{\tau})\left\lVert\sigma_{n}\right\rVert^{2}\tau$
$+(2\theta-1-\lambda)[L_{2}-L_{3}|X_{n+1}|^{2}-\left\lVert\sigma_{n+1}\right\rVert^{2}]\tau+\lambda L_{2}\tau+O_{n}.$
Taking the conditional expectation $\mathbb{E}[\cdot\mid\mathscr{F}_{n}]$ on both sides of the above inequality, noting the fact that $\mathbb{E}[O_{n}\mid\mathscr{F}_{n}]=0$, and discarding the terms $-2(1-\theta)L_{3}|X_{n}|^{2}$ and $-(2\theta-1-\lambda)L_{3}|X_{n+1}|^{2}$ under the assumption $\lambda+1-2\theta\leq 0$, $n\in\mathbb{N}$, we have
$\mathbb{E}[F_{n+1}+(2\theta-1-\lambda)\left\lVert\sigma_{n+1}\right\rVert^{2}\tau\mid\mathscr{F}_{n}]$
$\leq$ $N_{\tau}F_{n}+[1-B_{\tau}-2(1-\theta)]\left\lVert\sigma_{n}\right\rVert^{2}\tau+L_{2}\tau$
$\leq$ $N_{\tau}[F_{n}+(2\theta-1-\lambda)\left\lVert\sigma_{n}\right\rVert^{2}\tau]+L_{2}\tau,$
where in the last estimate, we have used the inequality
$1-B_{\tau}-2(1-\theta)<N_{\tau}(2\theta-1-\lambda),$
which is ensured by the estimate
$\frac{1+(1+\lambda)L_{3}\tau}{1+(1-\theta+\lambda)L_{3}\tau}>1.$
Therefore, (3.3) follows from the above estimate by subtracting the same deterministic term $L_{2}(1-\theta+\lambda)\tau$ from both sides.
Finally, we pointed out that the function $V_{\theta}:\mathbb{R}^{d}\to[0,\infty)$ defined by (3.2) satisfies $\lim_{x\to\infty}V_{\theta}(x)=\infty$ so that it is a Lyapunov function. Indeed,
$V_{\theta}(x)\geq$ $\left\lvert x\right\rvert^{2}-2(1-\theta+\lambda)\left\langle x,b(x)\right\rangle\tau+(2\theta-1-\lambda)\left\lVert\sigma(x)\right\rVert^{2}\tau$
$\geq$ $\left\lvert x\right\rvert^{2}-(1-\theta+\lambda)[L_{2}-L_{3}\left\lvert x\right\rvert^{2}-\left\lVert\sigma(x)\right\rVert^{2}]\tau+(2\theta-1-\lambda)\left\lVert\sigma(x)\right\rVert^{2}\tau$
$=$ $[1+(1-\theta+\lambda)L_{3}\tau]\left\lvert x\right\rvert^{2}+\theta\left\lVert\sigma(x)\right\rVert^{2}\tau-L_{2}(1-\theta+\lambda)\tau$
(3.5) $\geq$ $[1+(1-\theta+\lambda)L_{3}\tau]\left\lvert x\right\rvert^{2}-L_{2}(1-\theta+\lambda)\tau,$
which tends to $\infty$ as $x\to\infty$. ∎
Remark 3.1.
  1. (1)
    It follows from the estimation (3.3) that
    $\mathbb{E}V_{\theta}(X_{n})\leq\rho^{n}\mathbb{E}V_{\theta}(X_{0})+\frac{L_{2}\tau}{1-\rho}\leq\mathbb{E}V_{\theta}(X_{0})+\frac{L_{2}\tau}{1-\rho},\quad n\in\mathbb{N},$
    with $\rho=\frac{1+(1-\theta)L_{3}\tau}{1+(1-\theta+\lambda)L_{3}\tau}$, which shows the uniform moment stability of $(V_{\theta}(X_{n}))_{n\geq 1}$.
  2. (2)
    For the backward Euler scheme, a special case of (STM) with $\theta=1$, one can take the Lyapunov function (with $\lambda=1$) as
    $V_{1}(x)=\left\lvert x-b(x)\tau\right\rvert^{2},\quad x\in\mathbb{R}^{d},$
    which was derived in [28, Section 8.2] for dissipative problems.
When $\theta=1/2$, we have the following weak Lyapunov structure for (STM), which is sufficient, in combination with the irreducibility and strong Feller property in the next part, to derive the unique ergodicity, even without convergence rate, of the mid-point scheme, i.e., (STM) with $\theta=1/2$.
Corollary 3.1.
Let Assumption 2.1 and 2.2 hold. Then for $\tau\in(0,1)$, there exists a Lyapunov function $V_{1/2}:\mathbb{R}^{d}\to[0,\infty)$ defined by
(3.6) $V_{1/2}(x)=|x-\frac{1}{2}b(x)\tau|^{2},\quad x\in\mathbb{R}^{d},$
such that (STM) with $\theta=1/2$ satisfies
(3.7) $\mathbb{E}[V_{1/2}(X_{n+1})\mid\mathscr{F}_{n}]\leq V_{1/2}(X_{n})+L_{2}\tau-L_{3}|X_{n}|^{2},\quad n\in\mathbb{N}.$
Consequently, (2.6) holds with $c=L_{2}\tau+1$ and $C:=\{x\in\mathbb{R}^{d}:~{}x^{2}\leq\frac{L_{2}\tau+1}{L_{3}}\}$.
Proof.
As in the proof of Theorem 3.1, we choose $\theta=1/2$ and $\lambda=0$, so
$F_{n}=|X_{n}-\frac{1}{2}b_{n}\tau|^{2}+\frac{1}{2}L_{2}\tau,$
and
$F_{n+1}=F_{n}+2\left\langle X_{n},b_{n}\right\rangle\tau+\left\lVert\sigma_{n}\right\rVert^{2}\tau+O_{n},$
where $O_{n}=2\langle X_{n}+\frac{1}{2}b_{n}\tau,\sigma_{n}\delta_{n}W\rangle+\left\lvert\sigma_{n}\delta_{n}W\right\rvert^{2}-\left\lVert\sigma_{n}\right\rVert^{2}\tau$ satisfies $\mathbb{E}[O_{n}\mid\mathscr{F}_{n}]=0$, $n\in\mathbb{N}$.
By (2.3), we have
$F_{n+1}\leq$ $F_{n}+(L_{2}-L_{3}|X_{n}|^{2}-\left\lVert\sigma_{n}\right\rVert^{2})\tau+\left\lVert\sigma_{n}\right\rVert^{2}\tau+\lambda L_{2}\tau+O_{n}.$
Taking the conditional expectation $\mathbb{E}[\cdot\mid\mathscr{F}_{n}]$ on both sides of the above inequality and noting the fact that $\mathbb{E}[O_{n}\mid\mathscr{F}_{n}]=0$, we have
$\mathbb{E}[F_{n+1}\mid\mathscr{F}_{n}]\leq F_{n}+L_{2}\tau-L_{3}|X_{n}|^{2},$
and thus shows (3.7) with $V$ defined in (3.6).
Finally, the inequality (2.6) with $c=L_{2}\tau+1$ and $C:=\{x\in\mathbb{R}^{d}:~{}x^{2}\leq\frac{L_{2}\tau+1}{L_{3}}\}$ follows from (3.7) by direct calculations. ∎
3.2. Irreducibility and strong Feller property of STM
To derive the required probabilistic regularity for (STM), in this part, we denote by $P(x,A)$, $x\in\mathbb{R}^{d}$, $A\in\mathcal{B}(\mathbb{R}^{d})$, the transition kernel of the Markov chain $(X_{n})_{n\in\mathbb{N}}$ generated by (STM). Then
(3.8) $P(x,A)=\mathbb{P}(X_{n+1}\in A\mid X_{n}=x),\quad x\in\mathbb{R}^{d},~{}A\in\mathcal{B}(\mathbb{R}^{d}).$
We have the following irreducibility and existence of continuous density for (STM).
Proposition 3.1.
Let Assumption 2.1 and 2.3 hold. Then for any $\theta\in[1/2,1]$ and $\tau\in(0,1)$ with $L_{1}\theta\tau<2$, $(X_{n})_{n\in\mathbb{N}}$ is irreducible and strong Feller in $\mathbb{R}^{d}$. Consequently, there exists at most, if it exists, one invariant measure of $(X_{n})_{n\in\mathbb{N}}$ in $\mathbb{R}^{d}$.
Proof.
Let $x\in\mathbb{R}^{d}$, $A\in\mathcal{B}(\mathbb{R}^{d})$ be a non-empty open set, $\hat{b}$ be defined as in (2.2), and $\widetilde{b}(x):=x+(1-\theta)b(x)\tau$. By (3.8),
$P(x,A)$ $=\mathbb{P}(X_{n+1}\in A\mid X_{n}=x)$
$=\mathbb{P}(\hat{b}(x)=\widetilde{b}(x)+\sigma(x)\delta_{n}W\in\hat{b}(A))$
(3.9) $=\mu_{\widetilde{b}(x),\sigma(x)\sigma(x)^{\top}\tau}(\hat{b}(A)),$
as $\widetilde{b}(x)+\sigma(x)\delta_{n}W\sim\mathcal{N}(\widetilde{b}(x),\sigma(x)\sigma(x)^{\top}\tau)$, where $\mathcal{N}(a,b)$ and $\mu_{a,b}$ denote, respectively, the normal distribution and Gaussian measure in $\mathbb{R}^{d}$ with mean $a\in\mathbb{R}^{d}$ and variance operator $b\in\mathcal{L}(\mathbb{R}^{d})$. It suffices to show the irreducibility and strong Feller property of $(X_{n})_{n\in\mathbb{N}}$ in $\mathbb{R}^{d}$.
It follows from Assumption 2.3 that $\hat{b}$ is continuously differentiable. Then, by the inverse function theorem, $\hat{b}$ is an open map that maps open sets to open sets. Due to the non-degeneracy of $\sigma$ in Assumption 2.3, the Gaussian measure $\mu_{\widetilde{b}(x),\sigma(x)\sigma(x)^{\top}\tau}$ is non-degenerate, so that the open set $\hat{b}(A)$ has positive measure under $\mu_{\widetilde{b}(x),\sigma(x)\sigma(x)^{\top}\tau}$, which implies $P(x,A)>0$ and shows the irreducibility of $(X_{n})_{n\in\mathbb{N}}$ in $\mathbb{R}^{d}$.
Now fix a set $A\in\mathcal{B}(\mathbb{R}^{d})$. Define $f(x):=P(x,A)$, $x\in\mathbb{R}^{d}$. As $\widetilde{b}$ and $\sigma\sigma^{\top}$ are both continuous, by Lebesgue dominated convergence theorem, we have
$\lim_{y\to x}\mu_{\widetilde{b}(y),\sigma(y)\sigma(y)^{\top}\tau}(C)=\mu_{\widetilde{b}(x),\sigma(x)\sigma(x)^{\top}\tau}(C),\quad\forall~{}C\in\mathcal{B}(\mathbb{R}^{d}),$
from which we conclude $\lim_{y\to x}f(y)=f(x)$. Therefore, $f$ is continuous so that $(X_{n})_{n\in\mathbb{N}}$ is strong Feller in $\mathbb{R}^{d}$. ∎
3.3. Ergodicity of STM
In combination with the Lyapunov structures developed in Theorem 3.1 and Corollary 3.1 and minorization condition in Proposition 3.1, we have the following ergodicity of (STM).
Theorem 3.2.
Let Assumptions 2.1, 2.2, and 2.3 hold. Then the Markov chain $(X_{n})_{n\in\mathbb{N}}$ generated by (STM) is uniquely ergodic for any $\theta\in[1/2,1]$ and $\tau\in(0,1)$ with $L_{1}\theta\tau<2$.
Proof.
The Lyapunov conditions (3.3) and (3.7) in Theorem 3.1 and Corollary 3.1, imply the existence of an invariant measure for $(X_{n})_{n\in\mathbb{N}}$ in the cases $\theta\in(1/2,1]$ and $\theta=1/2$, respectively. In combination with the irreducibility and strong Feller property in Proposition 3.1, we conclude the uniqueness of the invariant measure for $(X_{n})_{n\in\mathbb{N}}$ with any $\theta\in[1/2,1]$. ∎
Under more regularity on the diffusions function $\sigma$, we can obtain the following geometric ergodicity of (STM) with $\theta\in(1/2,1]$.
Corollary 3.2.
In addition to the assumptions in Theorem 3.2, assume that the inverse $[\sigma\sigma^{\top}]^{-1}$ of $\sigma\sigma^{\top}$ are continuously differentiable. Then for any $\theta\in(1/2,1]$ and $\tau\in(0,1)$ with $L_{1}\theta\tau<2$, $(X_{n}^{x})_{n\in\mathbb{N}}$ generated by (STM) starting from $X_{0}=x$ admits a unique invariant measure $\pi_{\tau}^{\theta}$ which is geometric ergodic, i.e., there exists $r\in(0,1)$ and $\kappa\in(0,\infty)$ such that for all measurable $\phi$ with $|\phi|\leq V$,
$\left\lvert\mathbb{E}\phi(X_{n}^{x})-\pi_{\tau}^{\theta}(\phi)\right\rvert\leq\kappa r^{n}V(x),\quad\forall~{}x\in\mathbb{R}^{n}.$
Proof.
Due to [28, Lemma 2.3 and Theorem 2.5], in addition to the Lyapunov condition (3.3) and the irreducibility in Proposition 3.1, it suffices to show the existence of a jointly continuous density for the transition kernel $P(x,\cdot)$ of $(X_{n})$.
Indeed, for any $A\in\mathcal{B}(\mathbb{R}^{d})$, it follows from (3.2) that
$P(x,A)=$ $\int_{\hat{b}(A)}\frac{\exp\{-(y-\widetilde{b}(x))^{\top}[\sigma(x)\sigma(x)^{\top}\tau]^{-1}(y-\widetilde{b}(x))/2\}}{\sqrt{(2\pi)^{d}\det{[\sigma(x)\sigma(x)^{\top}\tau]}}}\mathrm{d}y$
$=$ $\int_{A}\frac{\exp\{-(\hat{b}(y)-\widetilde{b}(x))^{\top}[\sigma(x)\sigma(x)^{\top}\tau]^{-1}(\hat{b}(y)-\widetilde{b}(x))/2\}}{\sqrt{(2\pi)^{d}\det{[\sigma(x)\sigma(x)^{\top}\tau]}}}\left\lvert\det J_{\hat{b}}(y)\right\rvert\mathrm{d}y.$
Therefore, $P(x,dy)$ possesses a density $p(x,y)$ with respect to the Lebesgue measure in $\mathbb{R}^{d}$ given by
(3.10) $p(x,y):=\frac{\exp\{-(\hat{b}(y)-\widetilde{b}(x))^{\top}[\sigma(x)\sigma(x)^{\top}\tau]^{-1}(\hat{b}(y)-\widetilde{b}(x))/2\}}{\sqrt{(2\pi)^{d}\det{[\sigma(x)\sigma(x)^{\top}\tau]}}}\left\lvert\det J_{\hat{b}}(y)\right\rvert,~{}~{}y\in\mathbb{R}^{d}.$
For each $x_{0},y_{0}\in\mathbb{R}^{d}$, $p(x_{0},y)$ is continuously differentiable in $y$ in the neighborhood of $x_{0}$ (thus locally uniformly continuous) and $\lim_{x\to x_{0}}p(x,y_{0})=p(x_{0},y_{0})$ exists. Applying Moore–Osgood Theorem, we have
[…]
which shows the joint continuity of the density $p$ defined above. ∎
Remark 3.2.
It should be pointed out that the (one-step) irreducibility may be lost for the following Milstein-type of scheme:
$Y_{j+1}$ $=Y_{j}+(1-\theta)b(Y_{j})\tau+\theta b(Y_{j+1})\tau+\sigma(Y_{j})\delta_{j}W$
$\quad+\sigma^{\prime}(Y_{j})\sigma(Y_{j})\Big{[}\int_{t_{j}}^{t_{j+1}}(W(r)-W(t_{j})){\rm d}W_{r}\Big{]},\quad j\in\mathbb{N},$
under slightly more regularity and commutativity conditions on $\sigma$ as in [34]. Indeed, for $d=m=1$, it is clear that the above Milstein scheme is equivalent to
$Y_{j+1}$ $=Y_{j}+(1-\theta)b(Y_{j})\tau+\theta b(Y_{j+1})\tau+\sigma(Y_{j})\delta_{j}W+\frac{1}{2}\sigma^{\prime}(Y_{j})\sigma(Y_{j})(\delta_{j}W)^{2},$
$j\in\mathbb{N}$. If $\sigma^{\prime}$ is invertible, for any $x\in\mathbb{R}^{d}$, it is not difficult to show the existence of an open set $A\subset\mathbb{R}^{d}$ such that $P(x,A)=0$.
4. Application to Monotone SPDE
In this section, we shall apply a similar methodology as in the previous section to derive the unique ergodicity of the following SPDE:
(4.1) ${\rm d}X(t,\xi)=(\Delta X(t,\xi)+f(X(t,\xi))){\rm d}t+g(X(t,\xi)){\rm d}W(t,\xi),$
under (homogenous) Dirichlet boundary condition (DBC) $X(t,\xi)=0$, $(t,\xi)\in\mathbb{R}_{+}\times\partial\mathscr{O}$, with initial value condition $X(0,\xi)=X_{0}(\xi)$, $\xi\in\mathscr{O}$, where the physical domain $\mathscr{O}\subset\mathbb{R}^{d}$ ($d=1,2,3$) is a bounded open set with smooth boundary. Here, $f$ is assumed to be monotone-type with polynomial growth, and $g$ satisfies the usual Lipschitz condition in an infinite-dimensional setting; see Assumptions 4.1 and 4.2, respectively. It is clear that (4.1) includes the following SACE, arising from phase transition in materials science by stochastic perturbation, as a special case:
(4.2) ${\rm d}X(t,\xi)=\Delta X(t,\xi){\rm d}t+\epsilon^{-2}(X(t,\xi)-X(t,\xi)^{3}){\rm d}t+{\rm d}W(t,\xi),$
under DBC, where the positive index $\epsilon\ll 1$ is the interface thickness; see, e.g., [1], [2], [6], [10], [14], [26], [27], [30] and references therein.
4.1. Preliminaries
We first introduce some notations and main assumptions in the infinite-dimensional case.
[…]
Let ${\bf Q}$ be a self-adjoint and positive definite linear operator in $H$. Denote $H_{0}:={\bf Q}^{1/2}H$ and by $(\mathcal{L}_{2}^{0}:=HS(H_{0};H),\|\cdot\|_{\mathcal{L}_{2}^{0}})$ the space of Hilbert–Schmidt operators from $H_{0}$ to $H$. The driven process $W$ in Eq. (4.1) is an $H$-valued $\bf Q$-Wiener process on $(\Omega,\mathscr{F},\mathbb{F},\mathbb{P})$, which has the Karhunen–Loève expansion $W(t)=\sum_{k\in\mathbb{N}}\sqrt{q_{k}}g_{k}\beta_{k}(t)$, $t\geq 0$. Here $\{g_{k}\}_{k\in\mathbb{N}_{+}}$ are the eigenvectors of $\bf Q$ and form an orthonormal basis of $H$, with respect to eigenvalues $\{q_{k}\}_{k=1}^{\infty}$, and $\{\beta_{k}\}_{k\in\mathbb{N}_{+}}$ are mutually independent 1-D Brownain motions in $(\Omega,\mathscr{F},\mathbb{F},\mathbb{P})$; see, e.g., [23, Section 2.1] for more details. We mainly focus on trace-class noise, i.e., Q is a trace-class operator.
Our main conditions on the coefficients of (4.1) are the following two assumptions.
Assumption 4.1.
There exist scalars $K_{i}\in\mathbb{R}$, $i=1,2,3,4,5$, and $q\geq 1$ such that
(4.3) $(f(\xi)-f(\eta))(\xi-\eta)\leq K_{1}(\xi-\eta)^{2},\quad\xi,\eta\in\mathbb{R},$
(4.4) $f(\xi)\xi\leq K_{2}|\xi|^{2}+K_{3},\quad\xi\in\mathbb{R},$
(4.5) $|f^{\prime}(\xi)|\leq K_{4}|\xi|^{q-1}+K_{5},\quad\xi\in\mathbb{R}.$
Throughout, we assume that $q\geq 1$ when $d=1,2$ and $q\in[1,3]$ when $d=3$, so that the Sobolev embeddings $\dot{H}^{1}\subset L^{2q}(\mathscr{O})\subset H$ holds. Then we can define the Nemytskii operator $F:\dot{H}^{1}\rightarrow\dot{H}^{-1}$ associated with $f$ by
(4.6) $F(x)(\xi):=f(x(\xi)),\quad x\in\dot{H}^{1},\ \xi\in\mathscr{O}.$
It follows from the monotone condition (4.3) and the coercive condition (4.4) that the operator $F$ defined in (4.6) satisfies
(4.7) ${}_{1}\langle x-y,F(x)-F(y)\rangle_{-1}\leq K_{1}\|x-y\|^{2},$ $\quad x,y\in\dot{H}^{1},$
(4.8) ${}_{1}\langle x,F(x)\rangle_{-1}\leq K_{2}\|x\|^{2}+K_{3},$ $\quad x\in\dot{H}^{1}.$
The inequality (4.8), in combination with the Poincaré inequality that
(4.9) $\|\nabla x\|^{2}\geq\lambda_{1}\|x\|^{2},\quad x\in\dot{H}^{1},$
where $\lambda_{1}$ denotes the first eigenvalue of $-\Delta$ in $H$, implies that
(4.10) ${}_{1}\langle x,\Delta x+F(x)\rangle_{-1}$ $\leq-(\lambda_{1}-K_{2})\|x\|^{2}+K_{3},\quad x\in\dot{H}^{1}.$
Denote by $G:H\rightarrow\mathcal{L}_{2}^{0}$ the Nemytskii operator associated with $g$:
(4.11) $G(x)g_{k}(\xi):=g(x(\xi))g_{k}(\xi),\quad x\in H,~{}k\in\mathbb{N},~{}\xi\in\mathscr{O}.$
The following Lipschitz continuity and linear growth conditions are our main conditions on the diffusion operator $G$ defined in (4.11).
Assumption 4.2.
There exist positive constants $K_{6}$, $K_{7}$, and $K_{8}$ such that
(4.12) $\|G(x)-G(y)\|_{\mathcal{L}_{2}^{0}}^{2}\leq K_{6}\|x-y\|^{2},$ $\quad x,y\in H,$
(4.13) $\|G(x)\|^{2}_{\mathcal{L}_{2}^{0}}\leq K_{7}\|x\|^{2}+K_{8},$ $\quad x\in H.$
With these preliminaries, Eq. (4.1) is equivalent to the following infinite-dimensional stochastic evolution equation:
(4.14) ${\rm d}X(t)=(\Delta X(t)+F(X(t))){\rm d}t+G(X(t)){\rm d}W,\quad t\geq 0,$
where the initial datum $X(0)\in H$ is assumed to vanish on the boundary $\partial\mathscr{O}$ of the physical domain throughout the present paper.
We discretize Eq. (4.14) in time with a DIE scheme, which can be viewed as a special case of the STM (STM) with $\theta=1$ in infinite dimension, and the resulting temporal semi-discrete problem is to find an $\dot{H}^{1}$-valued Markov chain $\{X_{j}:\ j\in\mathbb{N}\}$ such that
(DIE) $X_{j+1}=X_{j}+\tau[\Delta X_{j+1}+F(X_{j+1})]+G(X_{j})\delta_{j}W,$
starting from $X_{0}\in H$, $j\in\mathbb{N}$. It had been widely studied; see, e.g., [11, 13, 24, 27], and references therein.
To derive the unique solvability of the DIE scheme (DIE), we define $\hat{F}:\dot{H}^{1}\to\dot{H}^{-1}$ by
(4.15) $\hat{F}(x)=({\rm Id}-\tau\Delta)x-\tau F(x),\quad x\in\dot{H}^{1}.$
Here and what follows, ${\rm Id}$ denotes the identity operator in various Hilbert spaces if there is no confusion. Then (DIE) becomes
$\hat{F}(X_{j+1})=X_{j}+G(X_{j})\delta_{j}W,\quad j\in\mathbb{N}.$
Lemma 4.1.
Let Assumptions 4.1 hold with $(K_{1}-\frac{\lambda_{1}}{\lambda_{1}+1})\tau<1$. Then $\hat{F}:\dot{H}^{1}\to\dot{H}^{-1}$ defined in (4.15) is bijective so that the DIE scheme (DIE) can be uniquely solved. Morevoer, $\hat{F}$ is an open map, i.e. for each open set $A\in\mathcal{B}(\dot{H}^{1})$, $\hat{F}(A)$ is also an open set in $\mathcal{B}(\dot{H}^{-1})$.
Proof.
Following from (4.7) with $K_{1}$ replaced by $\hat{K}\geq K_{1}$, which automatically holds, we have for all $x,y\in\dot{H}^{1}$,
${}_{1}\langle x-y,\hat{F}(x)-\hat{F}(y)\rangle_{-1}$ $=\left\lVert x-y\right\rVert^{2}+\tau\left\lVert\nabla(x-y)\right\rVert^{2}-\tau{}_{1}\langle x-y,F(x)-F(y)\rangle_{-1}$
$\geq(1-\hat{K}\tau)\left\lVert x-y\right\rVert^{2}+\tau\left\lVert\nabla(x-y)\right\rVert^{2}$
$=(1-\hat{K}\tau)\left\lVert x-y\right\rVert_{1}^{2}+(-1+\hat{K}\tau+\tau)\left\lVert\nabla(x-y)\right\rVert^{2}$
$\geq(1-\hat{K}\tau+\frac{\lambda_{1}}{\lambda_{1}+1}\tau)\left\lVert x-y\right\rVert_{1}^{2}+(-1+\hat{K}\tau)\left\lVert\nabla(x-y)\right\rVert^{2},$
where we have used the Poincaré inequality (4.9) in the last inequality. Note that $01-\hat{K}\tau+\frac{\lambda_{1}}{\lambda_{1}+1}\tau>$ 0 and $-1+\hat{K}\tau\geq 0$ hold if and only if
(4.16) $\frac{1}{\tau}\leq\hat{K}<\frac{1}{\tau}+\frac{\lambda_{1}}{\lambda_{1}+1}.$
As $(K_{1}-\frac{\lambda_{1}}{\lambda_{1}+1})\tau<1$, we can pick $\hat{K}:=\frac{1}{\tau}\vee K_{1}$ so that both $\hat{K}\geq K_{1}$ and (4.16) are satisfied. Hence,
(4.17) $_{1}\langle x-y,$ $\hat{F}(x)-\hat{F}(y)\rangle_{-1}\geq C_{0}\left\lVert x-y\right\rVert_{1}^{2},\quad\forall~{}x,y\in\dot{H}^{1},$
with $C_{0}:=1-\hat{K}\tau+\frac{\lambda_{1}}{\lambda_{1}+1}\tau>0$, which shows that strictly monotonicity and injectivity of $\hat{F}$.
On the other hand, $\hat{F}$ is coercive in the sense that $\lim_{\|x\|_{1}\to\infty}\frac{{}_{1}\langle x,\hat{F}(x)\rangle_{-1}}{\|x\|_{1}}=\infty$. Indeed, it follows from (4.17) that
${}_{1}\langle x,\hat{F}(x)\rangle_{-1}\geq C_{0}\left\lVert x\right\rVert_{1}^{2}+{}_{-1}\langle x,\hat{F}(0)\rangle_{1}\geq C_{0}\left\lVert x\right\rVert_{1}^{2}-\|\hat{F}(0)\|_{-1}\left\lVert x\right\rVert_{1},\quad\forall~{}x\in\dot{H}^{1},$
The above monotonicity and coercivity, in combination with the boundedness and continuity of $\hat{F}$, indicated in [25, Section 4], yield the bijectivity of $\hat{F}$ by Browder–Minty Theorem (see e.g. [32, Theorem 10.49]).
It remains to show that $\hat{F}$ is an open map. From the strict monotonicity (4.17) and the fact that
$\|\hat{F}(x)-\hat{F}(y)\|_{-1}\left\lVert x-y\right\rVert_{1}\geq{}_{1}\langle x-y,\hat{F}(x)-\hat{F}(y)\rangle_{-1},\quad\forall~{}x,y\in\dot{H}^{1},$
we obtain
$\|\hat{F}(x)-\hat{F}(y)\|_{-1}\geq C_{0}\|x-y\|_{1},\quad\forall~{}x,y\in\dot{H}^{1}.$
This shows that
(4.18) $B_{\dot{H}^{-1}}(\hat{F}(x),r)\subset\hat{F}(B_{\dot{H}^{1}}(x,r/C_{0})),\quad~{}\forall~{}r>0.$
Fix an open set $A\in\mathcal{B}(\dot{H}^{1})$. Our target, $\hat{F}(A)$ being open, means that for each point $x\in\dot{H}^{1}$, there exists an open ball $B_{\dot{H}^{-1}}(\hat{F}(x),r_{0})$ with $r_{0}>0$ such that $B_{\dot{H}^{-1}}(\hat{F}(x),r_{0})\subset\hat{F}(A)$. Due to the inclusion (4.18), such ball exists if there exists $B_{\dot{H}^{1}}(x,r_{0}/C)\subset A$. But this is guaranteed because $A$ is open. ∎
Remark 4.1.
We can use the same idea as above to show the $\hat{b}$ defined in (2.2) in previous sections is also an open map by its uniform monotonicity, thus relieving the assumption of continuous differentiability in Assumption 2.3 for the proof of Proposition 3.1. However, the latter assumption is not that restrictive in the finite-dimensional case and makes the proof straightforward with the help of the inverse function theorem. Also, it cannot be relieved in Corollary 3.2 when the joint continuity of the transition probability density is needed.
We also need the following non-degeneracy of $G$, which is equivalent to the invertibility of $G(x)G(x)^{\top}$ for each $x\in H$ recalling that ${\bf Q}^{1/2}$ is a positive definite operator in $H$. Here and after, denotes the corresponding adjoint operator.
Assumption 4.3.
For any $x\in H$, $G(x){\bf Q}^{1/2}[G(x){\bf Q}^{1/2}]^{\top}$ is a positive definite operator in $\mathcal{L}(H)$.
4.2. Lyapunov structure of DIE
We have the following Lyapunov structure for the DIE scheme; see [24, Lemma 3.1] for a similar uniform moments estimate.
Theorem 4.1.
Let Assumptions 4.1 and 4.2 hold with $K_{2}+K_{7}/2<\lambda_{1}$. Then for any $\tau\in(0,1)$ with $(K_{1}-\frac{\lambda_{1}}{\lambda_{1}+1})\tau<1$ and $\epsilon\in(0,\lambda_{1}-K_{2})$, there exists a Lyapunov function $V:\dot{H}^{1}\to[0,\infty)$ defined by
(4.19) $V(x)=\|x\|^{2}+\frac{2\epsilon\tau}{[1+2(\lambda_{1}-K_{2}-\epsilon)\tau]\lambda_{1}}\|\nabla x\|^{2},\quad x\in\dot{H}^{1},$
such that (2.7) holds for the drirt-implicit Euler scheme (DIE), with $\rho=\frac{1+K_{7}\tau}{1+2(\lambda_{1}-K_{2}-\epsilon)\tau}\in(0,1)$ and $\kappa=\frac{(2K_{3}+K_{8})\tau}{1+2(\lambda_{1}-K_{2}-\epsilon)\tau}>0$, i.e.,
(4.20) $\mathbb{E}[V(X_{n+1})\mid\mathscr{F}_{n}]\leq\frac{1+K_{7}\tau}{1+2(\lambda_{1}-K_{2}-\epsilon)\tau}V(X_{n})+\frac{(2K_{3}+K_{8})\tau}{1+2(\lambda_{1}-K_{2}-\epsilon)\tau},\quad n\in\mathbb{N}.$
Proof.
[…]
$2\langle x-y,x\rangle=\|x\|^{2}-\|y\|^{2}+\|x-y\|^{2},\quad x,y\in H,$
and integration by parts formula, we have
$\|X_{k+1}\|^{2}-\|X_{k}\|^{2}+\|X_{k+1}-X_{k}\|^{2}+\frac{2\varepsilon}{\lambda_{1}}\|\nabla X_{k+1}\|^{2}\tau$
$=-2(1-\frac{\varepsilon}{\lambda_{1}})\|\nabla X_{k+1}\|^{2}\tau+2~{}_{1}\langle X_{k+1},F_{k+1}\rangle_{-1}\tau$
(4.21) $\quad+2\langle X_{k+1}-X_{k},G_{k}\delta_{k}W\rangle+2\langle X_{k},G_{k}\delta_{k}W\rangle,$
for any positive $\varepsilon$ with $\varepsilon<\lambda_{1}$; here and in the rest of the paper, $\varepsilon$ denotes an arbitrarily small positive constant which would differ in each appearance. Using the estimate (4.8) and Cauchy–Schwarz inequality, we have
$\|X_{k+1}\|^{2}-\|X_{k}\|^{2}+\frac{2\varepsilon}{\lambda_{1}}\|\nabla X_{k+1}\|^{2}\tau$
$\leq-2(\lambda_{1}-K_{2}-\varepsilon)\|X_{k+1}\|^{2}\tau+\|G_{k}\delta_{k}W\|^{2}+2\langle X_{k},G_{k}\delta_{k}W\rangle+2K_{3}\tau.$
It follows that
$(1+2(\lambda_{1}-K_{2}-\varepsilon)\tau)\|X_{k+1}\|^{2}+\frac{2\varepsilon}{\lambda_{1}}\|\nabla X_{k+1}\|^{2}\tau$
$\leq\|X_{k}\|^{2}+\|G_{k}\delta_{k}W\|^{2}+2\langle X_{k},G_{k}\delta_{k}W\rangle+2K_{3}\tau.$
Taking the conditional expectation $\mathbb{E}[\cdot\mid\mathscr{F}_{n}]$ on both sides, noting the fact that both $X_{k}$ and $G_{k}$ are independent of $\delta_{k}W$, using Itô isometry and (4.13), we get
$(1+2(\lambda_{1}-K_{2}-\varepsilon)\tau)\mathbb{E}[\|X_{k+1}\|^{2}\mid\mathscr{F}_{n}]+\frac{2\varepsilon}{\lambda_{1}}\tau\mathbb{E}[\|\nabla X_{k+1}\|^{2}\mid\mathscr{F}_{n}]$
$\leq(1+K_{7}\tau)\|X_{k}\|^{2}+(2K_{3}+K_{8})\tau,$
from which we obtain (4.20).
Finally, we need to note from the compact embedding $\dot{H}^{1}\subset H$ that $V$ defined in (4.19) is indeed a Lyapunov function. ∎
4.3. Irreducibility and strong Feller property of DIE
In this part, we denote by $P(x,A)$, $x\in H$, $A\in\mathcal{B}(H)$, the transition kernel of the Markov chain $(X_{n})_{n\in\mathbb{N}}$ generated by (DIE):
(4.22) $P(x,A)=\mathbb{P}(X_{j+1}\in A\mid X_{j}=x),\quad x\in H,~{}A\in\mathcal{B}(H).$
We have the following irreducibility and strong Feller property of (DIE) in $H$.
Proposition 4.1.
Let Assumptions 4.1, 4.2, and 4.3 hold with $(K_{1}-\frac{\lambda_{1}}{\lambda_{1}+1})\tau<1$. Then $(X_{n})_{n\in\mathbb{N}}$ is irreducible and strong Feller in $H$. Consequently, there exists at most, if it exists, one invariant measure of $(X_{n})_{n\in\mathbb{N}}$ in $H$.
Proof.
Let $x\in H$. Firstly, we would like to extend the symmetric Gaussian measure $\mu_{0}:=\mu_{0,G(x){\bf Q^{1/2}}[G(x){\bf Q^{1/2}}]^{\top}\tau}$ in $H$ to another Gaussian measure $\mu_{1}$ on the larger Banach space $\dot{H}^{-1}$, with the same covariance operator. It suffices to show the irreducibility and strong Feller property of $(X_{n})_{n\in\mathbb{N}}$ in $H$.
Recall that by Kuratowski theorem (see e.g. [23, p.69]), we have $\mathcal{B}(H)=\mathcal{B}(\dot{H}^{-1})\cap H$ and $\mathcal{B}(\dot{H}^{1})=\mathcal{B}(H)\cap\dot{H}^{1}$. Then it is clear that $\mu_{1}$ defined by $\mu_{1}(A):=\mu_{0}(A\cap H)$, $A\in\mathcal{B}(\dot{H}^{-1})$, is a probability measure on $\dot{H}^{-1}$ and satisfies
$\mu_{1}(\phi\in B)=\mu_{0}(\{\phi\in B\}\cap H)$ $=\mu_{0}(\phi|_{H}\in B),\quad\forall~{}\phi\in(\dot{H}^{-1})^{*},~{}B\in\mathcal{B}(\mathbb{R}).$
As $\phi|_{H}\in H^{*}$ and $\mu_{0}$ is symmetric Gaussian in $H$, we conclude that $\mu_{1}$ is symmetric Gaussian on $\dot{H}^{-1}$ and maintains the covariance structure.
(i) For any non-empty open set $A\in\mathcal{B}(\dot{H}^{1})$, as in (3.2), we have
(4.23) $P(x,A)$ $=\mathbb{P}(X_{n+1}\in A\mid X_{n}=x)=\mu_{1}^{x}(\hat{F}(A)),$
[…]
Let $A\in\mathcal{B}(H)$ be a non-empty open set. It is evident by the Poincaré inequality (4.9) that there exists a non-empty open set $A_{1}\in\mathcal{B}(\dot{H}^{1})$ with $A_{1}\subset A$. Then we have $P(x,A)\geq P(x,A_{1})>0$, showing the irreducibility in $H$.
(ii) For any $A\in\mathcal{B}(\dot{H}^{1})$, as $\hat{F}$ and $G{\bf Q^{1/2}}[G{\bf Q^{1/2}}]^{\top}\tau$ are both continuous, by considering the characteristic function and Lebesgue dominated convergence theorem, we obtain
$\lim_{y\to x}P(y,A)=\lim_{y\to x}\mu_{1}^{y}(\hat{F}(A))=\mu_{1}^{x}(\hat{F}(A))=P(x,A).$
Now, let $A\in\mathcal{B}(H)$. Then $A$ can be decomposed as $A=A_{1}\cup A_{2}$ with $A_{1}:=A\cap\dot{H}^{1}\in\mathcal{B}(\dot{H}^{1})$ and $A_{2}:=A\cap(H\setminus\dot{H}^{1})$ being disjoint. It is clear that the DIE scheme (DIE) is almost surely $\dot{H}^{1}$-valued, by Lemma 4.1, so that $P(x,A_{2})=0$ for any $x\in H$. Then we have
$P(x,A)=P(x,A_{1})+P(x,A_{2})=P(x,A_{1}),\quad\forall~{}x\in H.$
The above two equalities imply
$\lim_{y\to x}P(y,A)=\lim_{y\to x}P(y,A_{1})=P(x,A_{1})=P(x,A),$
from which we conclude the continuity of $P(\cdot,A)$ in $H$ for any $A\in\mathcal{B}(H)$, indicating the strong Feller property of $(X_{n})_{n\in\mathbb{N}}$ in $H$. ∎
Theorem 4.2.
Let Assumptions 4.1, 4.2, and 4.3 hold with $K_{2}+K_{7}/2<\lambda_{1}$. Then the Markov chain $(X_{n})_{n\in\mathbb{N}}$ generated by the DIE scheme (DIE), is uniquely ergodic for any $\tau\in(0,1)$ with $(K_{1}-\frac{\lambda_{1}}{\lambda_{1}+1})\tau<1$.
Proof.
The Lyapunov condition (4.20) in Theorem 4.1, with $\rho=\frac{1+K_{7}\tau}{1+2(\lambda_{1}-K_{2}-\epsilon)\tau}\in(0,1)$ under the condition $K_{2}+K_{7}/2<\lambda_{1}$, implies the existence of an invariant measure for $(X_{n})_{n\in\mathbb{N}}$ generated by (DIE). In combination with the irreducibility and strong Feller property in Proposition 4.1, we conclude the uniqueness of the invariant measure for $(X_{n})_{n\in\mathbb{N}}$. ∎
4.4. Generalizations
In this part, we generalize the arguments and results in the previous three parts to Galerkin-based DIE and then to the SACE (4.2).
To introduce the fully discrete scheme, let $h\in(0,1)$, $\mathcal{T}_{h}$ be a regular family of partitions of $\mathscr{O}$ with maximal length $h$, and $V_{h}\subset\dot{H}^{1}$ be the space of continuous functions on $\bar{\mathscr{O}}$ which are piecewise linear over $\mathcal{T}_{h}$ and vanish on the boundary $\partial\mathscr{O}$. Denote by $\Delta_{h}:V_{h}\rightarrow V_{h}$ and $\mathcal{P}_{h}:\dot{H}^{-1}\rightarrow V_{h}$ be the discrete Laplacian and generalized orthogonal projection operators, respectively, defined by
(4.24) $\langle v^{h},\Delta_{h}x^{h}\rangle$ $=-\langle\nabla x^{h},\nabla v^{h}\rangle,\quad x^{h},v^{h}\in V_{h},$
(4.25) $\langle v^{h},\mathcal{P}_{h}z\rangle$ $=_{1}\langle v^{h},z\rangle_{-1},\quad z\in\dot{H}^{-1},\ v^{h}\in V_{h}.$
For $N\in\mathbb{N}_{+}$, denote by $V_{N}$ be the space spanned by the first $N$-eigenvectors of the Dirichlet Laplacian operator which vanish on $\partial\mathscr{O}$. Similarly, one can define the spectral Galerkin approximate Laplacian operator $\Delta_{N}:V_{N}\rightarrow V_{N}$ and generalized orthogonal projection operators $\mathcal{P}_{N}:\dot{H}^{-1}\rightarrow V_{N}$, respectively, as
$\langle\Delta_{N}u^{N},v^{N}\rangle$ $=-\langle\nabla u^{N},\nabla v^{N}\rangle,\quad u^{N},v^{N}\in V_{N},$
$\langle\mathcal{P}_{N}u,v^{N}\rangle$ $=_{1}\langle v^{N},u\rangle_{-1},\quad u\in V^{*},\ v^{N}\in V_{N}.$
Then the DIE Galerkin (DIEG) scheme of (4.14), or the Galerkin approximation of (DIE), is to find a $V^{h}$-valued discrete process $\{X^{h}_{j}:\ j\in\mathbb{N}\}$ such that
(DIEG) $X^{h}_{j+1}=X^{h}_{j}+\tau\Delta_{h}X^{h}_{j+1}+\tau\mathcal{P}_{h}F(X^{h}_{j+1})+\mathcal{P}_{h}G(X^{h}_{j})\delta_{j}W,$
$j\in\mathbb{N}$, starting from $X^{h}_{0}=\mathcal{P}_{h}X_{0}$, which had been widely studied; see, e.g., [13, 27]. One can also consider a spectral Galerkin version of the DIE scheme, where $\Delta_{h}$ and $\mathcal{P}_{h}$ in (DIEG) are replaced by $\Delta_{N}$ and $\mathcal{P}_{N}$, respectively.
Theorem 4.3.
Let Assumptions 4.1, 4.2, and 4.3 hold with $K_{2}<\lambda_{1}$. Then the Markov chain $(X_{n}^{h})_{n\in\mathbb{N}}$ generated by the DIEG scheme (DIEG) is uniquely ergodic for any $h\in(0,1)$ and $\tau\in(0,1)$ with $(K_{1}-\frac{\lambda_{1}}{\lambda_{1}+1})\tau<1$.
Proof.
At first, using the same arguments in Lemma 4.1 and the properties of the discrete Laplacian operator $\Delta_{h}$ and generalized orthogonal projection operator $\mathcal{P}_{h}$, defined in (4.24) and (4.25), respectively, we have the unique solvability of the DIEG scheme (DIEG) when $(K_{1}-\frac{\lambda_{1}}{\lambda_{1}+1})\tau<1$.
Secondly, the arguments in Theorem 4.1, in combination with the properties of $\Delta_{h}$ and $\mathcal{P}_{h}$, yields that $V:V_{h}\subset\dot{H}^{1}\to[0,\infty)$ defined in (4.19) is also a Lyapunov function of (DIEG) such that (4.20) holds with $X_{n}$ and $X_{n+1}$ replaced by $X_{n}^{h}$ and $X_{n+1}^{h}$, respectively.
Finally, as in the proof of Proposition 4.1, the irreducibility and strong Feller property of $(X_{n}^{h})_{n\in\mathbb{N}}$ can be derived by the Gaussian property of the driving ${\bf Q}$-Wiener process and the invertibility of $\mathcal{P}_{h}G{\bf Q}^{1/2}[\mathcal{P}_{h}G{\bf Q}^{1/2}]^{\top}$ followed from the non-degeneracy in Assumption 4.3. Therefore, we can conclude the unique ergodicity of (DIEG). ∎
Applying the above results in Theorem 4.2 and 4.3, we have the following unique ergodicity of DIE and Galerkin-based DIE (DIEG) applied to the SACE (4.2), which solves a question proposed in [11, Page 88].
Theorem 4.4.
For any $\epsilon>0$, $h\in(0,1)$ and $\tau\in(0,1)$ with $(\epsilon^{-2}-\frac{\lambda_{1}}{\lambda_{1}+1})\tau<1$, the numerical schemes (DIE) and (DIEG) for (4.2) are both uniquely ergodic. ∎
Proof.
We just need to check the conditions in Assumptions 4.1, 4.2, and 4.3 hold with $K_{2}<\lambda_{1}$ in the setting of (4.2) with $q=3$, $f(x)=\epsilon^{-2}(x-x^{3})$, $x\in\mathbb{R}$, and $G(x)={\rm Id}$ for all $x\in H$.
As $G(x)={\rm Id}$ for all $x\in H$, $G(x)G(x)^{\top}$ is invertible so that Assumption 4.3 holds. The validity of (4.3) and (4.5) in Assumption 4.1 and (4.12) and (4.13) in Assumption 4.2 in the setting of Eq. (4.2) were shown in [24, Section 4]: $K_{1}=\epsilon^{-2}$, $K_{4}=2\epsilon^{-2}$ and $K_{5}=\epsilon^{-2}$, $K_{6}=K_{7}=0$, $K_{8}={\rm Trace}({\bf Q})$, the trace of ${\bf Q}$. It remains to show (4.4) with $K_{2}<\lambda_{1}$ and some $K_{3}>0$. Indeed, by Young inequality,
$\epsilon^{-2}(x-x^{3})x=-\epsilon^{-2}x^{4}+\epsilon^{-2}x^{2}\leq-Cx^{2}+C_{\epsilon},$
for any positive $C$ and certain positive constant $C_{\epsilon}$. So one can take $K_{2}$ as any negative scalar and thus (4.4) with $K_{2}<\lambda_{1}$ and some $K_{3}>0$. ∎
5. Numerical Experiments
In this section, we perform two numerical experiments to verify our theoretical results Theorem 3.2 in Section 3 and Theorem 4.4 in Section 4, respectively.
The first numerical test is given to the SODE (SDE) with $b(x)=x-x^{3}$ and $\sigma(x)=\sqrt{x^{2}+1}$. Assumptions 2.1, 2.2, and 2.3 have been verified in Remark 2.1 with $L_{1}=2$, $L_{2}=3$, and $L_{3}=1$ and thus (STM) is uniquely ergodic for any $\theta\in[1/2,1]$ and $\tau\in(0,1)$ (so that $L_{1}\theta\tau<2$), according to Theorem 3.2. The proposed scheme, which is implicit, is numerically solved by utilizing scipy.optimize.fsolve, a wrapper around MINPACK’s hybrd and hybrj algorithms. In addition, we take $\tau=0.1$ and choose $\theta=1/2,3/4,1$ and initial data $X_{0}=-5,5,15$ to implement the numerical experiments.
It is clear from Figure 1 that the shapes of the empirical density functions plotted by kernel density estimation at $n=5000$ corresponding to $t=500$ are much more similar with the same $\theta$ and different initial data, which indicates the unique ergodicity of (STM) and thus verifies the theoretical result in Theorem 3.2. Indeed, Figure 1 indicates the strong mixing property (which yields the unique ergodicity) of (STM); this verifies the convergence result in Corollary 3.2 with $\theta\in(1/2,1]$. Moreover, the limiting measures for different $\theta$ are pretty close, which indicates the uniqueness of the limiting invariant measure for (STM) even with different $\theta$. Indeed, these invariant measures are all acceptable approximations of the limiting invariant measure, which is the unique invariant measure of Eq. (SDE).
Figure 1. Empirical density of (STM) for (SDE)
The second numerical test is given to the SACE (4.2) in $\mathscr{O}=(0,1)$ with $\epsilon=0.5$. By Theorem 4.4, (DIEG), as well as the spectral Galerkin discretization of (DIE), is uniquely ergodic for any $\tau\in(0,(4-\pi^{2}/(\pi^{2}+1))^{-1})$ (with $(\epsilon^{-2}-\frac{\lambda_{1}}{\lambda_{1}+1})\tau<1$). We take $\tau=0.05$ and $N=10$ (the dimension of the spectral Galerkin approximate space), choose $\theta=1/2,3/4,1$ and initial data $X_{0}(\xi)=0,\sin\pi\xi,\sum_{k=1}^{10}\sin k\pi\xi$, $\xi\in(0,1)$, and approximate the expectation by taking averaged value over $1,000$ paths to implement the numerical experiments. In addition, we simulate the time averages $\frac{1}{2,000}\sum_{n=1}^{2,000}\mathbb{E}[\phi(X_{n})]$ (up to $n=2,000$ corresponding to $t=100$) by
$\frac{1}{2,000,000}\sum_{n=1}^{2,000}\sum_{k=1}^{1,000}\phi(X_{n}^{k}),$
where $X_{n}^{k}$ denotes $n$-th iteration of $k$-th sample path and the test function $\phi$ are chosen to be $\phi(\cdot)=e^{-\|\cdot\|^{2}},\sin\|\cdot\|^{2},\|\cdot\|^{2}$, respectively.
Figure 2. Time averages of (DIEG) for (4.2)
From Figure 2, the time averages of (DIEG) with different initial data converge to the same ergodic limit, which verifies the theoretical result in Theorem 4.4.
6. Discussions
The theoretical loss of geometric ergodicity of (STM) with $\theta=1/2$ (in Corollary 3.2) is mainly because we can not show the stronger Lyapunov condition (3.3) at this stage. However, Figure 1 indicates that (STM) with $\theta=1/2$ is also strongly mixing. This motivates our conjecture that when $\theta=1/2$, or even for $\theta\in(0,1/2]$, (2.7) also holds with some $\rho\in(0,1)$ and $\kappa\geq 0$ so that it is also geometrically ergodic. The conjecture will be investigated in future research.