### v4

# Numerical Unique Ergodicity of Monotone SDEs driven by Nondegenerate Multiplicative Noise

Abstract.
We first establish the unique ergodicity of the stochastic theta method (STM) with $\theta\in[1/2,1]$ for monotone SODEs, without growth restriction on the coefficients, driven by nondegenerate multiplicative noise. The main ingredient of the arguments lies in constructing new Lyapunov functions involving the coefficients, the stepsize, and $\theta$ and deriving a minorization condition for the STM. We then generalize the arguments to the Galerkin-based full discretizations for a class of monotone SPDEs driven by infinite-dimensional nondegenerate multiplicative trace-class noise. Applying these results to the stochastic Allen–Cahn equation indicates that its Galerkin-based full discretizations are uniquely ergodic for any interface thickness. Numerical experiments verify our theoretical results.
Key words and phrases:
monotone stochastic differential equation, numerical invariant measure, numerical ergodicity, stochastic Allen–Cahn equation, Lyapunov structure
2010 Mathematics Subject Classification:
Primary 60H35; 60H15, 65M60
The first author is supported by the National Natural Science Foundation of China, No. 12101296, Basic and Applied Basic Research Foundation of Guangdong Province, No. 2024A1515012348, and Shenzhen Basic Research Special Project (Natural Science Foundation) Basic Research (General Project), No. JCYJ20220530112814033.
1. Introduction
The long-time behavior of Markov processes generated by stochastic differential equations (SDEs) is a natural and intriguing question and has been investigated in recent decades. As a significant long-time behavior, the ergodicity characterizes the case of temporal average coinciding with spatial average, which has a lot of applications in quantum mechanics, fluid dynamics, financial mathematics, and many other fields [12, 18]. The spatial average, i.e., the mean of a given test function for the invariant measure of the considered Markov process, also known as the ergodic limit, is desirable to compute in practical applications. One has to investigate a stochastic system over long time intervals, which is one of the main difficulties from the computational perspective. It is well-known that the explicit expression of the invariant measure for a stochastic nonlinear system is rarely available; exceptional examples are gradient Langevin systems driven by additive noise, e.g., [3, 21]. For this reason, it motivated and fascinated a lot of investigations in the recent decade for constructing numerical algorithms that can inherit the ergodicity of the original system and approximate the ergodic limit efficiently.
Much progress has been made in the design and analysis of numerical approximations of the desired ergodic limits for SDEs under a strong dissipative condition so that the Markov chains are contractive, see, e.g., [23, 35] and references therein for numerical ergodicity of backward Euler or Milstein schemes for dissipative SODEs, [22] for dissipative SODEs with Markovian switching, and [4, 5, 7, 8, 9, 11, 26] for approximating the invariant measures via temporal tamed, Galerkin-based linearity-implicit Euler or exponential Euler schemes, and high order integrator for parabolic SPDEs driven by additive noise. See also [29] and [14, 16] for numerical ergodicity of backward Euler scheme and its versions for monotone SODEs and spectral Galerkin approximation for 2-D stochastic Navier–Stokes equations, respectively, both driven by additive degenerate noise.
We note that most of the above works of literature focus on the numerical ergodicity of strong dissipative SDEs driven by additive noise; the numerical ergodicity for weak dissipative SDEs in the multiplicative noise case is more subtle. This question on the unique ergodicity of numerical approximations for monotone SDEs in finite and infinite-dimensional settings motivates the present study. Our main aim is to establish the unique ergodicity of the STM scheme (STM) with $\theta\in[1/2,1]$, including the numerical schemes studied in [17, 29, 32] for monotone SODEs, without growth restriction on the coefficients, driven by nondegenerate multiplicative noise.
It is not difficult to show that $V(\cdot)=\left\lvert\cdot\right\rvert^{2}$ is a natural Lyapunov function of the considered monotone (SODE). However, it was shown in [19] that the Euler–Maruyama scheme (i.e., (STM) with $\theta=0$) applied to Eq. (SODE) with superlinear growth coefficients would blow up in $p$-th moment for all $p\geq 2$. Therefore, $\left\lvert\cdot\right\rvert^{2}$ is not an appropriate Lyapunov function of this scheme in the setting of the present study, and we mainly focus on the case $\theta\in[1/2,1]$. By exploring the monotone structure of the coefficients and martingale property of the driven Wiener process, we construct several Lyapunov functions (see, e.g., (3.2) and (3.7) in Theorem 3.2 and Corollary 3.4, respectively) for the STM, which involve both the coefficients, the stepsize, and $\theta$. In combination with a minorization condition followed by deriving the irreducibility and the existence of a jointly continuous density for the transition kernel (see Proposition 3.5), we conclude the geometric ergodicity of the proposed STM scheme (STM).
Then we generalize our methodology to Galerkin-based fully discrete schemes (see (DIEG)), which have been studied in [11, 15, 20, 25, 28], for monotone SPDEs with polynomial growth coefficients driven by infinite-dimensional nondegenerate multiplicative trace-class noise (see Theorem 4.6). Applying these results to the stochastic Allen–Cahn equation (SACE) driven by nondegenerate noise indicates that its Galerkin-based full discretizations are uniquely ergodic, respectively, for any interface thickness (see Theorem 4.8).
The paper is organized as follows. In Section 2, we give the principal assumptions on monotone SODEs and recall the ergodic theory of Markov chains that will be used throughout. The STM’s Lyapunov structure and minorization condition are explored in Section 3. In Section 4, we generalize the arguments in Section 3 to monotone SPDEs, including the SACE. The theoretical results are validated by numerical experiments in Section 5. We include some discussions in the last section.
2. Assumptions on SODEs and Useful Facts from Markov Chain Theory
In this section, we present the required main assumptions and recall the general ergodic theory of Markov chains that will be used throughout the paper.
2.1. Main Assumptions on SODEs
Let us first consider the $d$-dimensional SODE
(SODE) $\mathrm{d}X(t)=b(X(t))\mathrm{d}t+\sigma(X(t))\mathrm{d}W(t),~{}t\geq 0,$
driven by an $\mathbb{R}^{m}$-valued Wiener process $W$ on a complete filtered probability space $(\Omega,\mathscr{F},\mathbb{F}:=(\mathscr{F}(t))_{t\geq 0},\mathbb{P})$, where $b:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ and $\sigma:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times m}$ are measurable functions.
Our main focus is on the invariant measure of the Markov chain generated by the following stochastic theta method (STM) with $\theta\in[0,1]$:
(STM) $X_{j+1}=X_{j}+(1-\theta)b(X_{j})\tau+\theta b(X_{j+1})\tau+\sigma(X_{j})\delta_{j}W,$
where $\tau\in(0,1)$ is a fixed step-size, $\delta_{j}W:=W(t_{j+1})-W(t_{j})$ with $t_{j}=j\tau$, $j\in\mathbb{N}$. When $\theta=0,1/2,1$, it is called the Euler–Maruyama scheme, the mid-point scheme, and the backward Euler method, respectively. We shall write $\mathscr{F}_{j}:=\mathscr{F}(t_{j})$ for notational convenience.
[…]
Assumption 2.1.
There exists a constant $L_{1}\in\mathbb{R}$ such that
(2.1) $2\left\langle b(x)-b(y),x-y\right\rangle+\left\lVert\sigma(x)-\sigma(y)\right\rVert^{2}\leq L_{1}\left\lvert x-y\right\rvert^{2},\quad\forall~{}x,y\in\mathbb{R}^{d}.$
Under Assumption (2.1), one can show the existence and uniqueness of the $(\mathscr{F}_{t})_{t\geq 0}$-adapted solution to (SODE), see, e.g., [24, Theorem 3.1.1]. The following lemma gives the analogous results for our numerical scheme (STM). Therefore, we shall always let this assumption hold to ensure the scheme we are talking about is a well-defined (homogenous) Markov chain.
Lemma 2.1.
Let Assumption 2.1 hold. Then STM scheme (STM) applied to Eq. (SODE) can be uniquely solved when $L_{1}\theta\tau<2$. Moreover, $(X_{n})_{n\in\mathbb{N}}$ is adapted to $(\mathscr{F}_{n})_{n\in\mathbb{N}}$ and enjoys homogenous Markov property.
Proof.
Define an auxiliary function $\hat{b}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ by
(2.2) $\hat{b}(x):=x-\theta b(x)\tau,\quad x\in\mathbb{R}^{d}.$
Our numerical scheme (STM) can be rewritten as
$\hat{b}(X_{j+1})=X_{j}+(1-\theta)b(X_{j}^{h})\tau+\sigma(X_{j})\delta_{j}W,\quad j\in\mathbb{N}.$
We want to show the invertibility of $\hat{b}$ through uniform monotonicity. Note that
$\langle x-y,\hat{b}(x)-\hat{b}(y)\rangle$ $=\left\langle x-y,x-y-(b(x)-b(y))\right\rangle\theta\tau$
$=\left\lVert x-y\right\rVert^{2}-\left\langle x-y,b(x)-b(y)\right\rangle\theta\tau$
$\geq(1-L_{1}\theta\tau/2)\left\lVert x-y\right\rVert^{2}.$
When $L_{1}\theta\tau<2$, we see that $\hat{b}$ defined in (2.2) is uniformly monotone and thus invertible (see, e.g., [34, Theorem C.2]), from which we can conclude the unique solvability of (STM).
As $X_{j+1}$ can be written as a function of $X_{j}$ and $\delta_{j}W=W(t_{j+1})-W(t_{j})$ (which is $\mathscr{F}_{j+1}$-measurable), one can easily show by induction that $(X_{n})_{n\in\mathbb{N}}$ is adapted to $(\mathscr{F}_{n})_{n\in\mathbb{N}}$. Finally, the homogenous Markov property is evident as $\delta_{j}W$ is stationary distributed and independent of $X_{j}$. ∎
Our second assumption is the following coupled coercive condition.
Assumption 2.2.
There exist two positive constants $L_{2}$ and $L_{3}$ such that
(2.3) $2\left\langle b(x),x\right\rangle+\left\lVert\sigma(x)\right\rVert^{2}\leq L_{2}-L_{3}\left\lvert x\right\rvert^{2},\quad\forall~{}x\in\mathbb{R}^{d}.$
We also need the following nondegeneracy and regularity conditions. Here and after, denotes the corresponding transpose of a matrix.
Assumption 2.3.
  1. (1)
    For each $x\in\mathbb{R}^{d}$, the matrix $\sigma(x)\sigma(x)^{\top}\in\mathbb{R}^{d\times d}$ is positive definite.
  2. (2)
    $b$ and the inverse $[\sigma\sigma^{\top}]^{-1}$ of $\sigma\sigma^{\top}$ are continuously differentiable.
Our assumptions are motivated by the below example.
Example 2.2.
Consider $d=m=1$ and
$b(x)=x-x^{3},\quad\sigma(x)=\sqrt{x^{2}+1},\quad x\in\mathbb{R}.$
Note that our choice of $\sigma$ is Lipschitz continuous with Lipschitz constant 1, so
$2\left\langle b(x)-b(y),x-y\right\rangle+\|\sigma(x)-\sigma(y)\|^{2}$
$=$ $2(x-y)^{2}(1-x^{2}-xy-y^{2})+|\sqrt{x^{2}+1}-\sqrt{y^{2}+1}|^{2}$
$\leq$ $\left\lvert x-y\right\rvert^{2}(3-2x^{2}-2xy-2y^{2})\leq 3\left\lvert x-y\right\rvert^{2},$
as $x^{2}+xy+y^{2}\geq 0$, for all $x,y\in\mathbb{R}$. This shows Assumption 2.1 holds with $L_{1}=3$. We now verify the second assumption:
$2\left\langle b(x),x\right\rangle+\|\sigma(x)\|^{2}=2x^{2}-2x^{4}+x^{2}+1\leq 3-x^{2}.$
This shows Assumption 2.2 with $L_{2}=3$, $L_{3}=1$. Finally, it is clear that Assumption 2.3 holds.
Remark 2.3.
The above example also shows that even if we have a negative coefficient $-L_{3}$ in the growth condition (2.3), we can still consider (2.1) with a positive coefficient $L_{1}$. These different dissipativity strengths would make the numerical ergodicity analysis very different; see recent work [23] for the strongly dissipative case.
Under assumptions very similar to us (Assumption 2.1, 2.2 and 2.3), Eq. (SODE) is ergodic, see, e.g., [33, 36] for the use of coupling methods based on a changing measure technique in combination of Girsanov theorem. In this paper, we focus on the numerical scheme, and such a Girsanov theorem does not exist in the discrete-time model. Instead, we will utilize the general ergodic theory of Markov chains.
2.2. Preliminaries on Ergodicity of Markov Chains
Let $(H,\|\cdot\|)$ (either $\mathbb{R}^{d}$ in Section 3 and $L^{2}$-space or its finite-dimensional approximate spaces in Section 4) be a separable Hilbert space equipped with Borel $\sigma$-algebra $\mathcal{B}(H)$. Denote by $\mathcal{N}(a,b)$ and $\mu_{a,b}$ the normal distribution and Gaussian measure in $H$, respectively, with mean $a\in H$ and variance operator $b\in\mathcal{L}(H)$. Let $(X_{n})_{n\in\mathbb{N}}$ be an $H$-valued Markov chain with transition kernel $P(x,B)$, $x\in H$, $B\in\mathcal{B}(H)$.
Recall that a set $C\in\mathcal{B}(H)$ is called an ($(m,\nu)$-)small set if there exists an integer $m>0$ and a non-trivial measure $\nu$ on $\mathcal{B}(H)$ such that
$P^{m}(x,B)\geq\nu(B),\quad\forall~{}x\in C,~{}B\in\mathcal{B}(H).$
$(X_{n})_{n\in\mathbb{N}}$ is said to satisfy the minorization condition (with $C$) if there exists some $\delta>0$, $C\in\mathcal{B}(H)$, and probability measure $\nu$ with $\nu(C)=1$ such that
(2.4) $P(x,B)\geq\delta\chi_{C}(x)\nu(B),\quad\forall~{}x\in H,~{}B\in\mathcal{B}(H).$
It is clear that if $(X_{n})_{n\in\mathbb{N}}$ satisfies the minorization condition (2.4) with $C$, then $C$ is a $(1,\delta\nu)$-small set. $(X_{n})_{n\in\mathbb{N}}$ is called (open set) irreducible if $P^{n}(x,A)>0$ for any $x\in H$, any non-empty open set $A$ in $H$ and some $n\in\mathbb{N}_{+}$; it is called strong Feller if $P(\cdot,A)$ is lower continuous for any Borel set $A\in\mathcal{B}(H)$. We refer to [30] for more details on Markov chains.
A probability measure $\mu$ in $H$ is called invariant for the Markov chain $(X_{n})_{n\in\mathbb{N}}$ or its transition kernel $P$, if
$\int_{H}P\phi(x)\mu({\rm d}x)=\mu(\phi):=\int_{H}\phi(x)\mu({\rm d}x),\quad\forall~{}\phi\in\mathcal{C}_{b}(H).$
This is equivalent to $\int_{H}P(x,A)\mu({\rm d}x)=\mu(A)$ for all $A\in\mathcal{B}(H).$ An invariant (probability) measure $\mu$ is called ergodic for $(X_{n})_{n\in\mathbb{N}}$ or $P$, if
(2.5) $\lim_{m\to\infty}\frac{1}{m}\sum_{k=0}^{m}P_{k}\phi(x)=\mu(\phi)\quad\text{in}~{}L^{2}(H;\mu),\quad\forall~{}\phi\in L^{2}(H;\mu).$
It is well-known that if $(X_{n})_{n\in\mathbb{N}}$ admits a unique invariant measure, then it is ergodic; in this case, we call it uniquely ergodic.
The unique ergodicity for a Markov chain is equivalent to the one-step Lyapunov condition
(2.6) $PV(x)-V(x)\leq-1+c\chi_{C}(x),\quad\forall~{}x\in H,$
for some $(m,\nu)$-small set $C$, constant $c$, and function $V:H\to[0,\infty)$ (called a Lyapunov function) which is finite $\nu$-almost everywhere (see, e.g., [30, Theorem 1.3.1]). A stronger but more adapted Lyapunov condition is that there exists $\rho\in(0,1)$ and $\kappa\geq 0$ such that
(2.7) $PV\leq\rho V+\kappa.$
Indeed, if (2.7) holds, then (2.6) holds with $C=\{V\leq(\kappa+1)/(1-\rho)\}$ and a large $c$. Under the stronger Lyapunov condition (2.7), one could obtain the geometric ergodicity if the minorization condition (2.4) holds for $C=\{x:V(x)\leq 2\kappa/(\gamma-\rho)\}$ with some $\gamma\in(\rho^{1/2},1)$ (see [29, Theorem 2.5]).
3. Geometric Ergodicity of STM for Monotone SODEs
Denote by $\mathcal{A}$ the infinitesimal generator of (SODE), i.e.,
$\mathcal{A}:=\sum_{i=1}^{d}b^{i}\frac{\partial}{\partial x_{i}}+\frac{1}{2}\sum_{i,j=1}^{d}(\sigma\sigma^{\top})_{ij}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}.$
Under Assumption 2.2, it is not difficult to show that
$\mathcal{A}\left\lvert X(t)\right\rvert^{2}=2\left\langle b(X(t)),X(t)\right\rangle+\left\lVert\sigma(X(t))\right\rVert^{2}\leq L_{2}-L_{3}\left\lvert X(t)\right\rvert^{2},\quad t\geq 0.$
Hence by utilizing Ito’s formula $V(\cdot)=\left\lvert\cdot\right\rvert^{2}$ is a natural Lyapunov function of (SODE); see e.g. [29, (2.2)] for more details. However, it was shown in [19] that for (STM) with $\theta=0$ applied to (SODE) with superlinear growth coefficients, $\lim_{n\to\infty}\mathbb{E}|X_{n}|^{p}=\infty$ for all $p\geq 2$ so that $V(x)=\left\lvert x\right\rvert^{p}$ is not an appropriate Lyapunov function of this scheme as the Lyapunov condition implies uniform moment stability (see Remark 3.3(i)).
In this section, we first construct a Lyapunov function and then derive the irreducibility and existence of continuous density for (STM) with $\theta\in[1/2,1]$. The case $\theta\in[0,1/2)$ is not covered.
3.1. Lyapunov structure
We begin with the following technical inequality, which will be helpful to explore the Lyapunov structure of (STM). It is a straightforward generalization of [32, Lemma 3.3]. Intuitively, the inequality says that we can throw away the term involving $\|\sigma(x)\|$ in the left-hand side by sacrificing a bit on the coefficients.
Lemma 3.1.
Let Assumption 2.2 hold. Then for any $\rho,\beta\in\mathbb{R}$ with $\rho\geq\beta\geq 0$, it holds with $C=\frac{1+\beta L_{3}}{1+\rho L_{3}}$ that for all $x\in\mathbb{R}^{d}$
(3.1) $\left\lvert x-\beta b(x)\right\rvert^{2}+\beta L_{2}+(C\rho-\beta)\|\sigma(x)\|^{2}\leq C(\left\lvert x-\rho b(x)\right\rvert^{2}+\rho L_{2})$
Proof.
Let $x\in\mathbb{R}^{d}$. By separating $x$ and $b(x)$, we have
$|x-$ $\beta b(x)|^{2}+\beta L_{2}+(C\rho-\beta)\|\sigma(x)\|^{2}-C(\left\lvert x-\rho b(x)\right\rvert^{2}+\rho L_{2})$
$=$ $(1-C)\left\lvert x\right\rvert^{2}+(C\rho-\beta)[2\left\langle x,b(x)\right\rangle-L_{2}+\|\sigma(x)\|^{2}]+(\beta^{2}-C\rho^{2})\left\lvert b(x)\right\rvert^{2}.$
Then it follows from (2.3) that the above equality can be bounded by
$[(1-C)-(C\rho-\beta)L_{3}]\left\lvert x\right\rvert^{2}+(\beta^{2}-C\rho^{2})\left\lvert b(x)\right\rvert^{2},$
from which we conclude (3.1) by noting that $1-C=(C\rho-\beta)L_{3}$ and $\beta^{2}-C\rho^{2}\leq 0$ with $C=\frac{1+\beta L_{3}}{1+\rho L_{3}}$. ∎
Now, we state and prove the following strong Lyapunov structure for (STM) with $\theta\in(1/2,1]$, which guarantees the existence of the invariant measure.
Theorem 3.2.
Let Assumption 2.1 and 2.2 hold and $\theta\in(1/2,1]$. For any $\tau\in(0,1)$ with $L_{1}\theta\tau<2$, and any constant $\lambda\in(0,2\theta-1]$, the function $V_{\theta}:\mathbb{R}^{d}\to[0,\infty)$ defined by
(3.2) $V_{\theta}(x):=\left\lvert x-(1-\theta+\lambda)b(x)\tau\right\rvert^{2}+(2\theta-1-\lambda)\left\lVert\sigma(x)\right\rVert^{2}\tau\quad x\in\mathbb{R}^{d},$
satisfies (2.7) with $\rho=\frac{1+(1-\theta)L_{3}\tau}{1+(1-\theta+\lambda)L_{3}\tau}\in(0,1)$ and $\kappa=L_{2}\tau>0$, i.e.,
(3.3) $\mathbb{E}[V_{\theta}(X_{n+1})\mid\mathscr{F}_{n}]\leq\frac{1+(1-\theta)L_{3}\tau}{1+(1-\theta+\lambda)L_{3}\tau}V_{\theta}(X_{n})+L_{2}\tau,\quad n\in\mathbb{N}.$
Moreover, $\lim_{x\to\infty}V_{\theta}(x)=\infty$, so that $V$ is a Lyapunov function.
Proof.
To simplify the notation in the following proof, we write as $b_{n}:=b(X_{n})$ and $\sigma_{n}:=\sigma(X_{n})$ and define
$F_{n}:=\left\lvert X_{n}-(1-\theta+\lambda)b_{n}\tau\right\rvert^{2}+L_{2}(1-\theta+\lambda)\tau.$
Remember that we aim to obtain a recursive relation for $(V(X_{n}))_{n\in\mathbb{N}}$. Firstly, we rearrange the terms to prepare for the substitution of $X_{n+1}$ by $X_{n}$ via their recursive relation:
$F_{n+1}=$ $\left\lvert X_{n+1}-\theta b_{n+1}\tau+(2\theta-1-\lambda)b_{n+1}\tau\right\rvert^{2}+L_{2}(1-\theta+\lambda)\tau$
$=$ $\left\lvert X_{n+1}-\theta b_{n+1}\tau\right\rvert^{2}+2(2\theta-1-\lambda)[\left\langle X_{n+1},b_{n+1}\right\rangle\tau-\theta\left\lvert b_{n+1}\right\rvert^{2}\tau^{2}]$
$\qquad+(2\theta-1-\lambda)^{2}\left\lvert b_{n+1}\right\rvert^{2}\tau^{2}+L_{2}(1-\theta+\lambda)\tau$
$=$ $\left\lvert X_{n+1}-\theta b_{n+1}\tau\right\rvert^{2}+2(2\theta-1-\lambda)\left\langle X_{n+1},b_{n+1}\right\rangle\tau$
$\qquad+(\lambda+1-2\theta)(\lambda+1)\left\lvert b_{n+1}\right\rvert^{2}\tau^{2}+L_{2}(1-\theta+\lambda)\tau.$
Substituting (STM) in the first term on the right-hand side of the above equality
$\left\lvert X_{n+1}-\theta b_{n+1}\tau\right\rvert^{2}=$ $\left\lvert X_{n}+(1-\theta)b_{n}\tau+\sigma_{n}\delta_{n}W\right\rvert^{2}$
$=$ $\left\lvert X_{n}+(1-\theta)b_{n}\tau\right\rvert^{2}+2\left\langle X_{n}+(1-\theta)b_{n}\tau,\sigma_{n}\delta_{n}W\right\rangle+\left\lvert\sigma_{n}\delta_{n}W\right\rvert^{2}$
$=$ $\left\lvert X_{n}-(1-\theta)b_{n}\tau\right\rvert^{2}+4(1-\theta)\left\langle X_{n},b_{n}\right\rangle\tau$
$\qquad+2\left\langle X_{n}+(1-\theta)b_{n}\tau,\sigma_{n}\delta_{n}W\right\rangle+\left\lvert\sigma_{n}\delta_{n}W\right\rvert^{2}.$
So that we obtain
(3.4) $\begin{split}F_{n+1}=&\left\lvert X_{n}-(1-\theta)b_{n}\tau\right\rvert^{2}+4(1-\theta)\left\langle X_{n},b_{n}\right\rangle\tau\\ &\qquad+2(2\theta-1-\lambda)\left\langle X_{n+1},b_{n+1}\right\rangle\tau\\ &\qquad+(\lambda+1-2\theta)(\lambda+1)\left\lvert b_{n+1}\right\rvert^{2}\tau^{2}+\left\lVert\sigma_{n}\right\rVert^{2}\tau\\ &\qquad+L_{2}(1-\theta+\lambda)\tau+O_{n},\end{split}$
where
$O_{n}=2\left\langle X_{n}+(1-\theta)b_{n}\tau,\sigma_{n}\delta_{n}W\right\rangle+\left\lvert\sigma_{n}\delta_{n}W\right\rvert^{2}-\left\lVert\sigma_{n}\right\rVert^{2}\tau.$
Note that $\delta_{n}W$ is independent of $\mathscr{F}_{n}$ so that by the normal distributed natural of $\delta_{n}W$, we see that $\mathbb{E}[O_{n}\mid\mathscr{F}_{n}]=0$.
Following (3.4) and Lemma 3.1 with $\rho=(1-\theta+\lambda)\tau$, $\beta=(1-\theta)\tau$, $C=N_{\tau}:=\frac{1+(1-\theta)L_{3}\tau}{1+(1-\theta+\lambda)L_{3}\tau}\in(0,1)$, and $C\rho-\beta=B_{\tau}:=\frac{\lambda\tau}{1+(1-\theta+\lambda)L_{3}\tau}$, and discarding the term $(\lambda+1-2\theta)(\lambda+1)\left\lvert b_{n+1}\right\rvert^{2}\tau^{2}$ under the assumption $10<\lambda\leq 2\theta-$ 1, we have
$F_{n+1}\leq$ $\left\lvert X_{n}-(1-\theta)b_{n}\tau\right\rvert^{2}+(1-\theta)L_{2}\tau+4(1-\theta)\left\langle X_{n},b_{n}\right\rangle\tau+\left\lVert\sigma_{n}\right\rVert^{2}\tau$
$\qquad+2(2\theta-1-\lambda)\left\langle X_{n+1},b_{n+1}\right\rangle\tau+L_{2}\lambda\tau+O_{n}$
$\leq$ $N_{\tau}F_{n}-B_{\tau}\|\sigma_{n}\|^{2}\tau+4(1-\theta)\left\langle X_{n},b_{n}\right\rangle\tau+\left\lVert\sigma_{n}\right\rVert^{2}\tau$
$\qquad+2(2\theta-1-\lambda)\left\langle X_{n+1},b_{n+1}\right\rangle\tau+L_{2}\lambda\tau+O_{n}$
$\leq$ $N_{\tau}F_{n}+2(1-\theta)(L_{2}-L_{3}|X_{n}|^{2}-\left\lVert\sigma_{n}\right\rVert^{2})\tau+(1-B_{\tau})\left\lVert\sigma_{n}\right\rVert^{2}\tau$
$\qquad+(2\theta-1-\lambda)[L_{2}-L_{3}|X_{n+1}|^{2}-\left\lVert\sigma_{n+1}\right\rVert^{2}]\tau+\lambda L_{2}\tau+O_{n}.$
Taking the conditional expectation $\mathbb{E}[\cdot\mid\mathscr{F}_{n}]$ on both sides of the above inequality, using the fact that $\mathbb{E}[O_{n}\mid\mathscr{F}_{n}]=0$, and discarding the terms $-2(1-\theta)L_{3}|X_{n}|^{2}$ and $-(2\theta-1-\lambda)L_{3}|X_{n+1}|^{2}$ under the assumption $\lambda+1-2\theta\leq 0$, $n\in\mathbb{N}$, we have
(3.5) $\mathbb{E}[F_{n+1}+$ $(2\theta-1-\lambda)\left\lVert\sigma_{n+1}\right\rVert^{2}\tau\mid\mathscr{F}_{n}]$
$\leq$ $N_{\tau}F_{n}+[1-B_{\tau}-2(1-\theta)]\left\lVert\sigma_{n}\right\rVert^{2}\tau+L_{2}\tau$
$\leq$ $N_{\tau}[F_{n}+(2\theta-1-\lambda)\left\lVert\sigma_{n}\right\rVert^{2}\tau]+L_{2}\tau,$
where in the last inequality, we have used
$1-B_{\tau}-2(1-\theta)<N_{\tau}(2\theta-1-\lambda),$
which is ensured by the estimate
$\frac{1+(1+\lambda)L_{3}\tau}{1+(1-\theta+\lambda)L_{3}\tau}>1.$
Therefore, the relation (3.3) follows from (3.5) by subtracting the same constant $L_{2}(1-\theta+\lambda)\tau$ from both sides.
Finally, we point out that the function $V_{\theta}:\mathbb{R}^{d}\to[0,\infty)$ defined by (3.2) satisfies $\lim_{x\to\infty}V_{\theta}(x)=\infty$ so that it is a Lyapunov function. Indeed,
$V_{\theta}(x)\geq$ $\left\lvert x\right\rvert^{2}-2(1-\theta+\lambda)\left\langle x,b(x)\right\rangle\tau+(2\theta-1-\lambda)\left\lVert\sigma(x)\right\rVert^{2}\tau$
$\geq$ $\left\lvert x\right\rvert^{2}-(1-\theta+\lambda)[L_{2}-L_{3}\left\lvert x\right\rvert^{2}-\left\lVert\sigma(x)\right\rVert^{2}]\tau+(2\theta-1-\lambda)\left\lVert\sigma(x)\right\rVert^{2}\tau$
$=$ $[1+(1-\theta+\lambda)L_{3}\tau]\left\lvert x\right\rvert^{2}+\theta\left\lVert\sigma(x)\right\rVert^{2}\tau-L_{2}(1-\theta+\lambda)\tau$
(3.6) $\geq$ $[1+(1-\theta+\lambda)L_{3}\tau]\left\lvert x\right\rvert^{2}-L_{2}(1-\theta+\lambda)\tau,$
which tends to $\infty$ as $x\to\infty$. ∎
Remark 3.3.
  1. (i)
    It follows from the estimation (3.3) that
    $\mathbb{E}V_{\theta}(X_{n})\leq\rho^{n}\mathbb{E}V_{\theta}(X_{0})+\frac{L_{2}\tau}{1-\rho}\leq\mathbb{E}V_{\theta}(X_{0})+\frac{L_{2}\tau}{1-\rho},\quad n\in\mathbb{N},$
    with $\rho=\frac{1+(1-\theta)L_{3}\tau}{1+(1-\theta+\lambda)L_{3}\tau}$, which shows the uniform moment stability of $(V_{\theta}(X_{n}))_{n\geq 1}$.
  2. (ii)
    For the backward Euler scheme, a special case of (STM) with $\theta=1$, one can take the Lyapunov function (with $\lambda=1$) as
    $V_{1}(x)=\left\lvert x-b(x)\tau\right\rvert^{2},\quad x\in\mathbb{R}^{d},$
    which coincides with the choice in [29, Section 8.2] for dissipative problems.
When $\theta=1/2$, we have the following weak Lyapunov structure for the mid-point scheme, which is sufficient, in combination with the irreducibility and strong Feller property in the next part, to derive the unique ergodicity.
Corollary 3.4.
Let Assumption 2.1 and 2.2 hold. For $\tau\in(0,1)$, the function $V_{1/2}:\mathbb{R}^{d}\to[0,\infty)$ defined by
(3.7) $V_{1/2}(x)=|x-\frac{1}{2}b(x)\tau|^{2},\quad x\in\mathbb{R}^{d},$
satisfies
(3.8) $\mathbb{E}[V_{1/2}(X_{n+1})\mid\mathscr{F}_{n}]\leq V_{1/2}(X_{n})+L_{2}\tau-L_{3}|X_{n}|^{2},\quad n\in\mathbb{N}.$
Consequently, (2.6) holds with $c=L_{2}\tau+1$ and $C:=\{x\in\mathbb{R}^{d}:~{}x^{2}\leq\frac{L_{2}\tau+1}{L_{3}}\}$. Moreover, $\lim_{x\to\infty}V_{1/2}(x)=\infty$ so that it is a Lyapunov function.
Proof.
We inherit the notations from the proof of Theorem 3.2 with $\theta=1/2$ and $\lambda=0$. Then, following the arguments there, we can define
$F_{n}=|X_{n}-\frac{1}{2}b_{n}\tau|^{2}+\frac{1}{2}L_{2}\tau,$
and have
$F_{n+1}=F_{n}+2\left\langle X_{n},b_{n}\right\rangle\tau+\left\lVert\sigma_{n}\right\rVert^{2}\tau+O_{n},$
where $O_{n}=2\langle X_{n}+\frac{1}{2}b_{n}\tau,\sigma_{n}\delta_{n}W\rangle+\left\lvert\sigma_{n}\delta_{n}W\right\rvert^{2}-\left\lVert\sigma_{n}\right\rVert^{2}\tau$ satisfies $\mathbb{E}[O_{n}\mid\mathscr{F}_{n}]=0$ for $n\in\mathbb{N}$.
By (2.3), we have
$F_{n+1}\leq$ $F_{n}+(L_{2}-L_{3}|X_{n}|^{2}-\left\lVert\sigma_{n}\right\rVert^{2})\tau+\left\lVert\sigma_{n}\right\rVert^{2}\tau+\lambda L_{2}\tau+O_{n}.$
Taking the conditional expectation $\mathbb{E}[\cdot\mid\mathscr{F}_{n}]$ on both sides of the above inequality and using $\mathbb{E}[O_{n}\mid\mathscr{F}_{n}]=0$, we have
$\mathbb{E}[F_{n+1}\mid\mathscr{F}_{n}]\leq F_{n}+L_{2}\tau-L_{3}|X_{n}|^{2},$
and thus shows (3.8) with $V$ defined in (3.7).
Finally, the inequality (2.6) with $c=L_{2}\tau+1$ and $C:=\{x\in\mathbb{R}^{d}:~{}x^{2}\leq\frac{L_{2}\tau+1}{L_{3}}\}$ follows from (3.8) by direct calculations. And the verification of $\lim_{x\to\infty}V_{1/2}(x)=\infty$ is the same as the previous case. ∎
3.2. Minorization condition and unique ergodicity of STM
To derive the required minorization condition for (STM), in this part, we denote by $P(x,A)$, $x\in\mathbb{R}^{d}$, $A\in\mathcal{B}(\mathbb{R}^{d})$, the transition kernel of the Markov chain $(X_{n})_{n\in\mathbb{N}}$ generated by (STM). Recall that for the homogenous Markov chain $(X_{n})_{n\in\mathbb{N}}$, we have for any $n\in\mathbb{N}$,
$P(x,A)=\mathbb{P}(X_{n+1}\in A\mid X_{n}=x),\quad x\in\mathbb{R}^{d},~{}A\in\mathcal{B}(\mathbb{R}^{d}).$
Let $x\in\mathbb{R}^{d}$, $A\in\mathcal{B}(\mathbb{R}^{d})$ be a non-empty open set, and set $\widetilde{b}(x):=x+(1-\theta)b(x)\tau$. Then, the above Markov property yields that
(3.9) $P(x,A)=\mathbb{P}(\widetilde{b}(x)+\sigma(x)\delta_{n}W\in\hat{b}(A))=\mu_{\widetilde{b}(x),\sigma(x)\sigma(x)^{\top}\tau}(\hat{b}(A)),$
as $\widetilde{b}(x)+\sigma(x)\delta_{n}W\sim\mathcal{N}(\widetilde{b}(x),\sigma(x)\sigma(x)^{\top}\tau)$, where $\hat{b}$ is defined in (2.2). Due to the nondegeneracy of $\sigma$ in Assumption 2.3(i), the Gaussian measure $\mu_{\widetilde{b}(x),\sigma(x)\sigma(x)^{\top}\tau}$ is nondegenerate. As we all know, in the finite-dimensional case, all non-degenerate Gaussian measures $\{\mu_{\widetilde{b}(x),\sigma(x)\sigma(x)^{\top}\tau}:x\in\mathbb{R}^{d}\}$ are equivalent to the underlying Lebesgue measure. Thus, $P$ is regular in the sense that all transition probabilities of the STM (STM) are equivalent.
Moreover, the following minorization condition holds for the STM scheme (STM).
Proposition 3.5.
Let Assumptions 2.1, 2.2, and 2.3 hold, and $\theta\in[1/2,1]$. Then for any $\tau\in(0,1)$ with $L_{1}\theta\tau<2$, $(X_{n}^{x})_{n\in\mathbb{N}}$ generated by (STM) satisfies the minorization condition (2.4) with a positive constant $\delta$, a small set $C\in\mathcal{B}(\mathbb{R}^{d})$, and a probability measure $\nu$ with $\nu(C)=1$.
Proof.
Due to [29, Lemma 2.3], it suffices to show the irreducibility and the existence of a jointly continuous density for the transition kernel $P(x,\cdot)$ of $(X_{n})$.
Let $A\in\mathcal{B}(\mathbb{R}^{d})$ be a non-empty open set. It follows from Assumption 2.3(ii) that $\hat{b}$ defined in (2.2) is continuously differentiable. Then, by the inverse function theorem, $\hat{b}$ is an open map that maps open sets to open sets so that the open set $\hat{b}(A)$ has a positive measure under the nondegenerate Gaussian measure $\mu_{\widetilde{b}(x),\sigma(x)\sigma(x)^{\top}\tau}$. This implies $P(x,A)>0$ and shows the irreducibility of $(X_{n})_{n\in\mathbb{N}}$ in $\mathbb{R}^{d}$.
To show the existence of a jointly continuous density for the transition kernel $P(x,\cdot)$, we set $A\in\mathcal{B}(\mathbb{R}^{d})$ and use (3.9) to derive
$P(x,A)=$ $\int_{\hat{b}(A)}\frac{\exp\{-(y-\widetilde{b}(x))^{\top}[\sigma(x)\sigma(x)^{\top}\tau]^{-1}(y-\widetilde{b}(x))/2\}}{\sqrt{(2\pi)^{d}\det{[\sigma(x)\sigma(x)^{\top}\tau]}}}\mathrm{d}y$
$=$ $\int_{A}\frac{\exp\{-(\hat{b}(y)-\widetilde{b}(x))^{\top}[\sigma(x)\sigma(x)^{\top}\tau]^{-1}(\hat{b}(y)-\widetilde{b}(x))/2\}}{\sqrt{(2\pi)^{d}\det{[\sigma(x)\sigma(x)^{\top}\tau]}}}\left\lvert\det J_{\hat{b}}(y)\right\rvert\mathrm{d}y.$
Therefore, $P(x,dy)$ possesses a density $p(x,y)$ with respect to the Lebesgue measure in $\mathbb{R}^{d}$ given by
(3.10) $p(x,y):=\frac{\exp\{-(\hat{b}(y)-\widetilde{b}(x))^{\top}[\sigma(x)\sigma(x)^{\top}\tau]^{-1}(\hat{b}(y)-\widetilde{b}(x))/2\}}{\sqrt{(2\pi)^{d}\det{[\sigma(x)\sigma(x)^{\top}\tau]}}}\left\lvert\det J_{\hat{b}}(y)\right\rvert,~{}~{}y\in\mathbb{R}^{d}.$
For each $x_{0},y_{0}\in\mathbb{R}^{d}$, $p(x_{0},y)$ is continuously differentiable in $y$ in the neighborhood of $x_{0}$ (thus locally uniformly continuous) and $\lim_{x\to x_{0}}p(x,y_{0})=p(x_{0},y_{0})$ exists. Once the uniform continuity on one variable is obtained, we can interchange to the order of limit and able to claim
[…]
which shows the joint continuity of the density $p$ defined above. ∎
Remark 3.6.
It should be pointed out that the (one-step) irreducibility may be lost for the following Milstein-type of scheme:
$Y_{j+1}$ $=Y_{j}+(1-\theta)b(Y_{j})\tau+\theta b(Y_{j+1})\tau+\sigma(Y_{j})\delta_{j}W$
$\quad+\sigma^{\prime}(Y_{j})\sigma(Y_{j})\Big{[}\int_{t_{j}}^{t_{j+1}}(W(r)-W(t_{j})){\rm d}W_{r}\Big{]},\quad j\in\mathbb{N},$
under slightly more regularity and commutativity conditions on $\sigma$ as in [35]. Indeed, for $d=m=1$, it is clear that the above Milstein scheme is equivalent to
$Y_{j+1}$ $=Y_{j}+(1-\theta)b(Y_{j})\tau+\theta b(Y_{j+1})\tau+\sigma(Y_{j})\delta_{j}W+\frac{1}{2}\sigma^{\prime}(Y_{j})\sigma(Y_{j})(\delta_{j}W)^{2},$
$j\in\mathbb{N}$. If $\sigma^{\prime}$ is invertible, for any $x\in\mathbb{R}^{d}$, it is not difficult to show the existence of an open set $A\subset\mathbb{R}^{d}$ such that $P(x,A)=0$.
In combination with the Lyapunov structures developed in Theorem 3.2 and Corollary 3.4 and the minorization condition derived in Proposition 3.5, we have the following uniform ergodicity of the STM scheme (STM) with $\theta=1/2$ and geometric ergodicity of the STM scheme (STM) with $\theta\in(1/2,1]$.
Theorem 3.7.
Let Assumptions 2.1, 2.2, and 2.3 hold. Then for any $\tau\in(0,1)$ with $L_{1}\theta\tau<2$, the STM scheme (STM) is uniquely ergodic with $\theta=1/2$, and it admits a unique invariant measure $\pi_{\tau}^{\theta}$ which is geometric ergodic with $\theta\in(1/2,1]$, i.e., there exists $r\in(0,1)$ and $\kappa\in(0,\infty)$ such that for all measurable $\phi$ with $|\phi|\leq V$,
$\left\lvert\mathbb{E}\phi(X_{n}^{x})-\pi_{\tau}^{\theta}(\phi)\right\rvert\leq\kappa r^{n}V(x),\quad\forall~{}x\in\mathbb{R}^{n}.$
4. Unique Ergodicity of Galerkin-based Full Discretizations for Monotone SPDEs
In this section, we shall apply a similar methodology as in the previous section to derive the unique ergodicity of Galerkin-based full discretizations for the following SPDE:
(SPDE) ${\rm d}X(t,\xi)=(\Delta X(t,\xi)+f(X(t,\xi))){\rm d}t+g(X(t,\xi)){\rm d}W(t,\xi),$
under (homogenous) Dirichlet boundary condition (DBC) $X(t,\xi)=0$, $(t,\xi)\in\mathbb{R}_{+}\times\partial\mathscr{O}$, with initial value condition $X(0,\xi)=X_{0}(\xi)$, $\xi\in\mathscr{O}$, where the physical domain $\mathscr{O}\subset\mathbb{R}^{d}$ ($d=1,2,3$) is a bounded open set with smooth boundary. Here, $f$ is assumed to be monotone-type with polynomial growth, and $g$ satisfies the usual Lipschitz condition in an infinite-dimensional setting; see Assumptions 4.1 and 4.2, respectively. It is clear that (SPDE) includes the following stochastic Allen–Cahn equation (SACE), arising from phase transition in materials science by stochastic perturbation, as a special case:
(SACE) ${\rm d}X(t,\xi)=\Delta X(t,\xi){\rm d}t+\epsilon^{-2}(X(t,\xi)-X(t,\xi)^{3}){\rm d}t+{\rm d}W(t,\xi),$
under DBC, where the positive index $\epsilon\ll 1$ is the interface thickness; see, e.g., [1], [2], [6], [10], [15], [27], [28], [31] and references therein.
4.1. Preliminaries and Galerkin-based Full Discretizations
We first introduce some notations and main assumptions in the infinite-dimensional case.
Denote by $\|\cdot\|_{H}$ and $\langle\cdot,\cdot\rangle_{H}$ the inner product and norm, respectively, in $H:=L^{2}(\mathscr{O})$. We usually omit the subscript $H$ if there is no confusion about the notations used for SODEs in Section 2. For $\theta=-1$ or $1$, we use $(\dot{H}^{\theta}=\dot{H}^{\theta}(\mathscr{O}),\|\cdot\|_{\theta})$ to denote the usual Sobolev interpolation spaces, respectively; the dual between $\dot{H}^{1}$ and $\dot{H}^{-1}$ are denoted by ,${}_{1}\langle\cdot,\cdot\rangle_{-1}$.
Let ${\bf Q}$ be a self-adjoint and positive definite linear operator on $H$. Denote $H_{0}:={\bf Q}^{1/2}H$ and by $(\mathcal{L}_{2}^{0}:=HS(H_{0};H),\|\cdot\|_{\mathcal{L}_{2}^{0}})$ the space of Hilbert–Schmidt operators from $H_{0}$ to $H$. The driven process $W$ in Eq. (SPDE) is an $H$-valued $\bf Q$-Wiener process on $(\Omega,\mathscr{F},\mathbb{F},\mathbb{P})$, which has the Karhunen–Loève expansion $W(t)=\sum_{k\in\mathbb{N}}\sqrt{q_{k}}g_{k}\beta_{k}(t)$, $t\geq 0$. Here $\{g_{k}\}_{k\in\mathbb{N}_{+}}$ are the eigenvectors of $\bf Q$ and form an orthonormal basis of $H$, with respect to eigenvalues $\{q_{k}\}_{k=1}^{\infty}$, and $\{\beta_{k}\}_{k\in\mathbb{N}_{+}}$ are mutually independent 1-D Brownain motions in $(\Omega,\mathscr{F},\mathbb{F},\mathbb{P})$; see, e.g., [24, Section 2.1] for more details. We mainly focus on trace-class noise, i.e., $\bf Q$ is a trace-class operator or equivalently, ${\rm Tr}({\bf Q}):=\sum_{k=1}^{\infty}q_{k}<\infty$.
Our main conditions on the coefficients of (SPDE) are the following two assumptions.
Assumption 4.1.
There exist scalars $K_{i}\in\mathbb{R}$, $i=1,2,3,4,5$, and $q\geq 1$ such that
(4.1) $(f(\xi)-f(\eta))(\xi-\eta)\leq K_{1}(\xi-\eta)^{2},\quad\xi,\eta\in\mathbb{R},$
(4.2) $f(\xi)\xi\leq K_{2}|\xi|^{2}+K_{3},\quad\xi\in\mathbb{R},$
(4.3) $|f^{\prime}(\xi)|\leq K_{4}|\xi|^{q-1}+K_{5},\quad\xi\in\mathbb{R}.$
Throughout, we assume that $q\geq 1$ when $d=1,2$ and $q\in[1,3]$ when $d=3$, so that the Sobolev embeddings $\dot{H}^{1}\subset L^{2q}(\mathscr{O})\subset H$ holds. Then we can define the Nemytskii operator $F:\dot{H}^{1}\rightarrow\dot{H}^{-1}$ associated with $f$ by
(4.4) $F(x)(\xi):=f(x(\xi)),\quad x\in\dot{H}^{1},\ \xi\in\mathscr{O}.$
It follows from the monotone condition (4.1) and the coercive condition (4.2) that the operator $F$ defined in (4.4) satisfies
(4.5) ${}_{1}\langle x-y,F(x)-F(y)\rangle_{-1}\leq K_{1}\|x-y\|^{2},$ $\quad x,y\in\dot{H}^{1},$
(4.6) ${}_{1}\langle x,F(x)\rangle_{-1}\leq K_{2}\|x\|^{2}+K_{3},$ $\quad x\in\dot{H}^{1}.$
The inequality (4.6), in combination with the Poincaré inequality that
(4.7) $\|\nabla x\|^{2}\geq\lambda_{1}\|x\|^{2},\quad x\in\dot{H}^{1},$
where $\lambda_{1}$ denotes the first eigenvalue of $-\Delta$ in $H$, implies that
(4.8) ${}_{1}\langle x,\Delta x+F(x)\rangle_{-1}$ $\leq-(\lambda_{1}-K_{2})\|x\|^{2}+K_{3},\quad x\in\dot{H}^{1}.$
Denote by $G:H\rightarrow\mathcal{L}_{2}^{0}$ the Nemytskii operator associated with $g$:
(4.9) $G(x)g_{k}(\xi):=g(x(\xi))g_{k}(\xi),\quad x\in H,~{}k\in\mathbb{N},~{}\xi\in\mathscr{O}.$
The following Lipschitz continuity and linear growth conditions are our main conditions on the diffusion operator $G$ defined in (4.9).
Assumption 4.2.
There exist positive constants $K_{6}$, $K_{7}$, and $K_{8}$ such that
(4.10) $\|G(x)-G(y)\|_{\mathcal{L}_{2}^{0}}^{2}\leq K_{6}\|x-y\|^{2},$ $\quad x,y\in H,$
(4.11) $\|G(x)\|^{2}_{\mathcal{L}_{2}^{0}}\leq K_{7}\|x\|^{2}+K_{8},$ $\quad x\in H.$
With these preliminaries, (SPDE) is equivalent to the following infinite-dimensional stochastic evolution equation:
(SEE) ${\rm d}X(t)=(\Delta X(t)+F(X(t))){\rm d}t+G(X(t)){\rm d}W,\quad t\geq 0,$
where the initial datum $X(0)\in H$ is assumed to vanish on the boundary $\partial\mathscr{O}$ of the physical domain throughout the present paper.
We also need to assume the nondegeneracy of $G$ in the following sense.
Assumption 4.3.
For any $x\in H$, $G(x)G(x)^{*}$ is a positive definite operator in $\mathcal{L}(H)$, i.e.,
$\langle G(x)G(x)^{*}h,h\rangle>0,\quad\mbox{for all non-zero $h\in H$}.$
Remark 4.1.
  1. (i)
    Together with the assumption that $\bf Q$ is self-adjoint and positive definite we made at the beginning of Section 4, Assumption 4.3 is equivalent to the positive definiteness of $[G(x){\bf Q}^{1/2}][G(x){\bf Q}^{1/2}]^{*}$ for any $x\in H$. Indeed, we observe that if the operator is of the form “$AA^{*}$”, then $\langle AA^{*}u,u\rangle>0$ if and only if $A^{*}u\neq 0$, since $\langle AA^{*}u,u\rangle=\langle A^{*}u,A^{*}u\rangle=\|A^{*}u\|^{2}$. Hence we have
    • $G(x)G(x)^{*}$ is positive definite if and only if $G^{*}u\neq 0$ for all non-zero $u\in H$;
    • $[G(x){\bf Q}^{1/2}][G(x){\bf Q}^{1/2}]^{*}$ is positive definite if and only if $[G(x){\bf Q}^{1/2}]^{*}u\neq 0$ for all non-zero $u\in H$; and moreover,
    • $\bf Q$ is positive definite if and only if ${\bf Q}^{1/2}u\neq 0$ for all non-zero $u\in H$ (recall that if $\bf Q$ is self-adjoint and positive definite, ${\bf Q}^{1/2}$ is also self-adjoint and positive definite).
    Let $u$ be a non-zero element of $H$. It is now clear that ${\bf Q}^{1/2}G(x)^{*}u\neq 0$ if and only if $G(x)^{*}u\neq 0$ by the positive definiteness of $\bf Q$. This shows the equivalence between the two conditions.
  2. (ii)
    (Readers may first read the notations we introduced below and then come back to the remark) Under Assumption 4.3, the finite-dimensional projection of the operator is also nondegenerate: for any $x\in H$,
    $[\mathcal{P}_{h}G(x){\bf Q}^{1/2}][\mathcal{P}_{h}G(x){\bf Q}^{1/2}]^{*}$
    is a positive definite operator in $\mathcal{L}(V_{h})$. To see this, it is enough to see that $\mathcal{P}_{h}^{*}u\neq 0$ for non-zero $u\in V_{h}$. Indeed, if $\mathcal{P}_{h}^{*}u=0$ for $u\in V_{h}$, then ${}_{1}\langle e_{k},\mathcal{P}_{h}^{*}u\rangle_{-1}=\langle\mathcal{P}_{h}e_{k},u\rangle=\langle e_{k},u\rangle=0$ for all $\{e_{k}\}_{k=1}^{K}$ the orthonormal basis of $V_{h}$. This implies $u=0$.
To introduce the Galerkin-based fully discrete scheme, let $h\in(0,1)$, $\mathcal{T}_{h}$ be a regular family of partitions of $\mathscr{O}$ with maximal length $h$, and $V_{h}\subset\dot{H}^{1}$ be the space of continuous functions on $\bar{\mathscr{O}}$ which are piecewise linear over $\mathcal{T}_{h}$ and vanish on the boundary $\partial\mathscr{O}$. Denote by $\Delta_{h}:V_{h}\rightarrow V_{h}$ and $\mathcal{P}_{h}:\dot{H}^{-1}\rightarrow V_{h}$ be the discrete Laplacian and generalized orthogonal projection operators, respectively, defined by
(4.12) $\langle v^{h},\Delta_{h}x^{h}\rangle$ $=-\langle\nabla x^{h},\nabla v^{h}\rangle,\quad x^{h},v^{h}\in V_{h},$
(4.13) $\langle v^{h},\mathcal{P}_{h}z\rangle$ $=_{1}\langle v^{h},z\rangle_{-1},\quad z\in\dot{H}^{-1},\ v^{h}\in V_{h}.$
We discretize Eq. (SEE) in time with a Drift-Implicit Euler (DIE) scheme, which can be viewed as a particular case of the STM (STM) with $\theta=1$ in infinite dimension, and in space with a Galerkin approximation. Then the resulting fully discrete, DIE Galerkin (DIEG) scheme of (SEE) is to find a $V^{h}$-valued discrete process $\{X^{h}_{j}:\ j\in\mathbb{N}\}$ such that
(DIEG) $X^{h}_{j+1}=X^{h}_{j}+\tau\Delta_{h}X^{h}_{j+1}+\tau\mathcal{P}_{h}F(X^{h}_{j+1})+\mathcal{P}_{h}G(X^{h}_{j})\delta_{j}W,$
$j\in\mathbb{N}$, starting from $X^{h}_{0}=\mathcal{P}_{h}X_{0}$, which had been widely studied; see, e.g., [13, 28].
Remark 4.2.
One can also consider a spectral Galerkin version of the DIEG scheme (DIEG) such that all results in this section will be valid, where $\Delta_{h}$ and $\mathcal{P}_{h}$ in (DIEG) are replaced by spectral Galerkin approximate Laplacian operator $\Delta_{N}:V_{N}\rightarrow V_{N}$ and the generalized orthogonal projection operators $\mathcal{P}_{N}:\dot{H}^{-1}\rightarrow V_{N}$, respectively:
$\langle\Delta_{N}u^{N},v^{N}\rangle$ $=-\langle\nabla u^{N},\nabla v^{N}\rangle,\quad u^{N},v^{N}\in V_{N},$
$\langle\mathcal{P}_{N}u,v^{N}\rangle$ $=_{1}\langle v^{N},u\rangle_{-1},\quad u\in V^{*},\ v^{N}\in V_{N},$
where $V_{N}$ denotes the space spanned by the first $N$-eigenvectors of the Dirichlet Laplacian operator which vanish on $\partial\mathscr{O}$ for $N\in\mathbb{N}_{+}$.
To derive the unique solvability of the DIEG scheme (DIEG), we define $\hat{F}:V_{h}\to V_{h}$ by
(4.14) $\hat{F}(x):=({\rm Id}-\tau\Delta_{h})x-\tau\mathcal{P}_{h}F(x),\quad x\in V_{h}.$
Here and what follows, ${\rm Id}$ denotes the identity operator in various Hilbert spaces if there is no confusion. Then (DIEG) becomes
$\hat{F}(X_{j+1}^{h})=X_{j}^{h}+\mathcal{P}_{h}G(X_{j}^{h})\delta_{j}W,\quad j\in\mathbb{N}.$
Lemma 4.3.
Let Assumptions 4.1 hold with $(K_{1}-\lambda_{1})\tau<1$. Then $\hat{F}:V_{h}\to V_{h}$ defined in (4.14) is bijective so that the DIE scheme (DIEG) can be uniquely solved and $(X_{j}^{h})$ is a $V_{h}$-valued homogenous Markov chain.
Morevoer, $\hat{F}$ is an open map, i.e. for each open set $A\in\mathcal{B}(\dot{V}_{h})$, $\hat{F}(A)$ is also an open set in $\mathcal{B}(V_{h})$.
Proof.
We have for all $x,y\in V_{h}$,
$\langle x-y,\hat{F}(x)-\hat{F}(y)\rangle$ $=\left\lVert x-y\right\rVert^{2}+\tau\left\lVert\nabla(x-y)\right\rVert^{2}-\tau{}_{1}\langle x-y,F(x)-F(y)\rangle_{-1}$
$\geq(1-K_{1}\tau)\left\lVert x-y\right\rVert^{2}+\tau\left\lVert\nabla(x-y)\right\rVert^{2}$
$\geq(1-K_{1}\tau+\lambda_{1}\tau)\left\lVert x-y\right\rVert^{2}$
where we have used the Poincaré inequality (4.7) in the last inequality. This shows the bijectivity of $\hat{F}$ (since $V_{h}$ is finite-dimensional, monotone uniform fixed point theorem holds). The homogenous Markov property is clear as the SODE case; see the proof of Lemma 2.1.
It remains to show that $\hat{F}$ is an open map. From the uniform monotonicity and Cauchy–Schwarz inequality, we obtain
(4.15) $\|\hat{F}(x)-\hat{F}(y)\|\geq(1-K_{1}\tau+\lambda_{1}\tau)\|x-y\|,\quad\forall~{}x,y\in V_{h}.$
As we explained below, $\hat{F}$ is an open map. From (4.15), we see that
(4.16) $B(\hat{F}(x),r)\subset\hat{F}(B(x,r/C_{0})),\quad~{}\forall~{}r>0,$
where $C_{0}:=1-K_{1}\tau+\lambda_{1}\tau>0$. Fix an open set $A\in\mathcal{B}(H)$. Our target, $\hat{F}(A)$ being open, means that for each point $x\in V_{h}$, there exists an open ball $B(\hat{F}(x),r_{0})$ with $r_{0}>0$ such that $B(\hat{F}(x),r_{0})\subset\hat{F}(A)$. Due to the inclusion (4.16), such ball exists if $B(x,r_{0}/C_{0})\subset A$ exists. But this is guaranteed because $A$ is open. ∎
Remark 4.4.
We can use the same idea as above to show the $\hat{b}$ defined in (2.2) in previous sections is also an open map by its uniform monotonicity, thus relieving the assumption of continuous differentiability in Assumption 2.3 for the proof of regularity (see the discussion above Proposition 3.5) of transition kernels of (STM). However, it cannot be relieved in Proposition 3.5 when the joint continuity of the transition probability density is needed.
4.2. Lyapunov structure of DIEG
We have the following Lyapunov structure for the DIEG scheme (DIEG); see [26, Lemma 3.1] for a similar uniform moments estimate for the temporal semi-discretization DIE scheme. Whether there exists a similar Lyapunov structure as that of (DIEG) for general $\theta\in[1/2,1)$ is unknown.
Theorem 4.5.
Let Assumptions 4.1 and 4.2 hold with $K_{2}+K_{7}/2<\lambda_{1}$. Then for any $\tau\in(0,1)$ with $(K_{1}-\lambda_{1})\tau<1$ and $\epsilon\in(0,\lambda_{1}-K_{2})$, there exists a Lyapunov function $V:V_{h}\to[0,\infty)$ defined by
(4.17) $V(x)=\|x\|^{2}+\frac{2\epsilon\tau}{[1+2(\lambda_{1}-K_{2}-\epsilon)\tau]\lambda_{1}}\|\nabla x\|^{2},\quad x\in V_{h},$
such that (2.7) holds for the DIEG scheme (DIEG), with $\rho=\frac{1+K_{7}\tau}{1+2(\lambda_{1}-K_{2}-\epsilon)\tau}\in(0,1)$ and $\kappa=\frac{(2K_{3}+K_{8})\tau}{1+2(\lambda_{1}-K_{2}-\epsilon)\tau}>0$, i.e.,
(4.18) $\mathbb{E}[V(X_{n+1}^{h})\mid\mathscr{F}_{n}]\leq\frac{1+K_{7}\tau}{1+2(\lambda_{1}-K_{2}-\epsilon)\tau}V(X_{n}^{h})+\frac{(2K_{3}+K_{8})\tau}{1+2(\lambda_{1}-K_{2}-\epsilon)\tau},\quad n\in\mathbb{N}.$
Proof.
[…]
$2\langle x-y,x\rangle=\|x\|^{2}-\|y\|^{2}+\|x-y\|^{2},\quad x,y\in H,$
and integration by parts formula, we have
$\|X_{k+1}^{h}\|^{2}-\|X_{k}^{h}\|^{2}+\|X_{k+1}^{h}-X_{k}^{h}\|^{2}+\frac{2\varepsilon}{\lambda_{1}}\|\nabla X_{k+1}^{h}\|^{2}\tau$
$=-2(1-\frac{\varepsilon}{\lambda_{1}})\|\nabla X_{k+1}^{h}\|^{2}\tau+2~{}_{1}\langle X_{k+1}^{h},F_{k+1}\rangle_{-1}\tau$
(4.19) $\quad+2\langle X_{k+1}^{h}-X_{k}^{h},G_{k}\delta_{k}W\rangle+2\langle X_{k}^{h},G_{k}\delta_{k}W\rangle,$
for any positive $\varepsilon$ with $\varepsilon<\lambda_{1}$; here and in the rest of the paper, $\varepsilon$ denotes an arbitrarily small positive constant which would differ in each appearance. Using the estimate (4.6) and Cauchy–Schwarz inequality, we have
$\|X_{k+1}^{h}\|^{2}-\|X_{k}^{h}\|^{2}+\frac{2\varepsilon}{\lambda_{1}}\|\nabla X_{k+1}^{h}\|^{2}\tau$
$\leq-2(\lambda_{1}-K_{2}-\varepsilon)\|X_{k+1}^{h}\|^{2}\tau+\|G_{k}\delta_{k}W\|^{2}+2\langle X_{k}^{h},G_{k}\delta_{k}W\rangle+2K_{3}\tau.$
It follows that
$(1+2(\lambda_{1}-K_{2}-\varepsilon)\tau)\|X_{k+1}^{h}\|^{2}+\frac{2\varepsilon}{\lambda_{1}}\|\nabla X_{k+1}^{h}\|^{2}\tau$
$\leq\|X_{k}^{h}\|^{2}+\|G_{k}\delta_{k}W\|^{2}+2\langle X_{k}^{h},G_{k}\delta_{k}W\rangle+2K_{3}\tau.$
Taking the conditional expectation $\mathbb{E}[\cdot\mid\mathscr{F}_{n}]$ on both sides, noting the fact that both $X_{k}^{h}$ and $G_{k}$ are independent of $\delta_{k}W$, using Itô isometry and (4.11), we get
$(1+2(\lambda_{1}-K_{2}-\varepsilon)\tau)\mathbb{E}[\|X_{k+1}^{h}\|^{2}\mid\mathscr{F}_{n}]+\frac{2\varepsilon}{\lambda_{1}}\tau\mathbb{E}[\|\nabla X_{k+1}^{h}\|^{2}\mid\mathscr{F}_{n}]$
$\leq(1+K_{7}\tau)\|X_{k}^{h}\|^{2}+(2K_{3}+K_{8})\tau,$
from which we obtain (4.18). ∎
Theorem 4.6.
Let Assumptions 4.1, 4.2, and 4.3 hold with $K_{2}<\lambda_{1}$. Then the DIEG scheme (DIEG) is uniquely ergodic for any $h\in(0,1)$ and $\tau\in(0,1)$ with $(K_{1}-\lambda_{1})\tau<1$.
Proof.
According to the Lyapunov structure derived in Theorem 4.5, it suffices to show the regularity of the transition kernel $P^{h}$ associated with the DIEG scheme (DIEG) defined as
$P^{h}(x,A)=\mathbb{P}(X_{n+1}^{h}\in A\mid X_{n}^{h}=x),\quad x\in V_{h},~{}A\in\mathcal{B}(V_{h}).$
By the Markov property of (DIEG) (similar to what we did in (3.9)) and calculations of mean and covariance, we have
(4.20) $P^{h}(x,A)=\mu_{x,[\mathcal{P}_{h}G(x){{\bf Q}^{1/2}}][\mathcal{P}_{h}G(x){{\bf Q}^{1/2}}]^{*}\tau}(\hat{F}(A)),\quad x\in V_{h},~{}A\in\mathcal{B}(V_{h}),$
[…]
Remark 4.7.
One can show that the estimate (4.18) holds for the temporal DIE scheme with the same function $V$ defined in (4.17), which is indeed a Lyapunov function in $H$ by the compact embedding $\dot{H}^{1}\subset H$. However, the regularity or the strong Feller property in the infinite-dimensional case is unknown, so one cannot conclude the unique ergodicity of the temporal semi-discrete DIE scheme.
Applying the above result in Theorem 4.6, we have the following unique ergodicity of the DIEG scheme (DIEG) applied to the SACE (SACE).
Theorem 4.8.
For any $\epsilon>0$, $h\in(0,1)$ and $\tau\in(0,1)$ with $(\epsilon^{-2}-\lambda_{1})\tau<1$, the numerical schemes (DIEG) and (DIEG) for (SACE) are both uniquely ergodic. ∎
Proof.
We just need to check the conditions in Assumptions 4.1, 4.2, and 4.3 hold with $K_{2}<\lambda_{1}$ in the setting of (SACE) with $q=3$, $f(x)=\epsilon^{-2}(x-x^{3})$, $x\in\mathbb{R}$, and $G(x)={\rm Id}$ for all $x\in H$.
As $G(x)={\rm Id}$ for all $x\in H$, $G(x)G(x)^{*}={\rm Id}$ is positive definite so that Assumption 4.3 holds. The validity of (4.1) and (4.3) in Assumption 4.1 and (4.10) and (4.11) in Assumption 4.2 in the setting of Eq. (SACE) were shown in [26, Section 4]: $K_{1}=\epsilon^{-2}$, $K_{4}=2\epsilon^{-2}$ and $K_{5}=\epsilon^{-2}$, $K_{6}=K_{7}=0$, $K_{8}={\rm Trace}({\bf Q})$, the trace of ${\bf Q}$. It remains to show (4.2) with $K_{2}<\lambda_{1}$ and some $K_{3}>0$. Indeed, by Young inequality,
$\epsilon^{-2}(x-x^{3})x=-\epsilon^{-2}x^{4}+\epsilon^{-2}x^{2}\leq-Cx^{2}+C_{\epsilon},$
for any positive $C$ and certain positive constant $C_{\epsilon}$. So one can take $K_{2}$ as any negative scalar and thus (4.2) with $K_{2}<\lambda_{1}$ and some $K_{3}>0$. ∎
5. Numerical Experiments
In this section, we perform two numerical experiments to verify our theoretical results Theorem 3.7 in Section 3 and Theorem 4.8 in Section 4, respectively.
5.1. Experiments on STM
The first numerical test is given to the SODE (SODE) with $b(x)=x-x^{3}$ and $\sigma(x)=\sqrt{x^{2}+1}$. Assumptions 2.1, 2.2, and 2.3 have been verified in Example 2.2 with $L_{1}=3$, $L_{2}=3$, and $L_{3}=1$ and thus (STM) is uniquely ergodic for any $\theta\in[1/2,1]$ and $\tau\in(0,1)$ (so that $L_{1}\theta\tau<2$), according to Theorem 3.7. The proposed scheme, which is implicit, is numerically solved by utilizing scipy.optimize.fsolve, a wrapper around MINPACK’s hybrd and hybrj algorithms. In addition, we take $\tau=0.1$ and choose $\theta=1/2,3/4,1$ and initial data $X_{0}=-5,5,15$ to implement the numerical experiments.
It is clear from Figure 1 that the shapes of the empirical density functions plotted by kernel density estimation at $n=5000$ corresponding to $t=500$ are much more similar with the same $\theta$ and different initial data, which indicates the unique ergodicity of (STM) and thus verifies the theoretical result in Theorem 3.7. Indeed, Figure 1 indicates the strong mixing property (which yields the unique ergodicity) of (STM); this verifies the convergence result in Theorem 3.7 with $\theta\in(1/2,1]$. Moreover, the limiting measures for different $\theta$ are pretty close, which indicates the uniqueness of the limiting invariant measure for (STM) even with different $\theta$. Indeed, these invariant measures are all acceptable approximations of the limiting invariant measure, which is the unique invariant measure of Eq. (SODE).
Figure 1. Empirical density of (STM) for (SODE)
5.2. Experiments on DIEG
The second numerical test is given to the SACE (SACE) in $\mathscr{O}=(0,1)$ with $\epsilon=0.5$. By Theorem 4.8, (DIEG), as well as the spectral Galerkin discretization of (DIEG), is uniquely ergodic for any $\tau\in(0,(4-\pi^{2}/(\pi^{2}+1))^{-1})$ (with $(\epsilon^{-2}-\frac{\lambda_{1}}{\lambda_{1}+1})\tau<1$). We take $\tau=0.05$ and $N=10$ (the dimension of the spectral Galerkin approximate space), choose $\theta=1/2,3/4,1$ and initial data $X_{0}(\xi)=0,\sin\pi\xi,\sum_{k=1}^{10}\sin k\pi\xi$, $\xi\in(0,1)$, and approximate the expectation by taking averaged value over $1,000$ paths to implement the numerical experiments. In addition, we simulate the time averages $\frac{1}{2,000}\sum_{n=1}^{2,000}\mathbb{E}[\phi(X_{n})]$ (up to $n=2,000$ corresponding to $t=100$) by
$\frac{1}{2,000,000}\sum_{n=1}^{2,000}\sum_{k=1}^{1,000}\phi(X_{n}^{k}),$
where $X_{n}^{k}$ denotes $n$-th iteration of $k$-th sample path and the test function $\phi$ are chosen to be $\phi(\cdot)=e^{-\|\cdot\|^{2}},\sin\|\cdot\|^{2},\|\cdot\|^{2}$, respectively.
Figure 2. Time averages of (DIEG) for (SACE)
From Figure 2, the time averages of (DIEG) with different initial data converge to the same ergodic limit, which verifies the theoretical result in Theorem 4.8.
6. Discussions
The theoretical loss of geometric ergodicity of (STM) with $\theta=1/2$ (in Theorem 3.7) is mainly because we can not show the stronger Lyapunov condition (3.3) at this stage. However, Figure 1 indicates that (STM) with $\theta=1/2$ is also strongly mixing. This motivates our conjecture that when $\theta=1/2$, or even for $\theta\in(0,1/2]$, (2.7) also holds with some $\rho\in(0,1)$ and $\kappa\geq 0$ so that it is also geometrically ergodic. The conjecture will be investigated in future research.